### v1

# Learn, Check, Test – Security Testing Using Automata Learning and Model Checking

Abstract
Cyber-physical systems are part of industrial systems and critical infrastructure. Therefore, they should be examined in a comprehensive manner to verify their correctness and security. At the same time, the complexity of such systems demands such examinations to be systematic and, if possible, automated for efficiency and accuracy. A method that can be useful in this context is model checking. However, this requires a model that faithfully represents the behavior of the examined system. Obtaining such a model is not trivial, as many of these systems can be examined only in black box settings due to, e.g., long supply chains or secrecy. We therefore utilize active black box learning techniques to infer behavioral models in the form of Mealy machines of such systems and translate them into a form that can be evaluated using a model checker. To this end, we will investigate a cyber-physical systems as a black box using its external communication interface. We first annotate the model with propositions by mapping context information from the respective protocol to the model using Context-based Proposition Maps (CPMs). We gain annotated Mealy machines that resemble Kripke structures. We then formally define a template, which we use to transfer the structures in to a format to be processed by a model checker. We further define generic security properties based on basic security requirements (authentication, confidentiality, privilege levels, and key validity). Due to the used CPMs, we can instantiate these properties with a meaningful context to check a specific protocol, which makes the approach flexible and scalable. Furthermore, the gained model can be easily altered to introduce non-deterministic behavior (like timeouts) or faults and examined if the properties still hold under these different conditions. Lastly, we demonstrate the versatility of the approach by providing case studies of very different systems (a passport and an automotive control unit), speaking different communication protocols (NFC and UDS), checked with the same tool chain and the same security properties.
keywords:
Automata Learning , Model Checking , Cybersecurity , NFC , UDS , Rebeca , Formal Methods
journal: Computers and Security
1 Introduction
Cyber-physical systems in all domains of everyday life become more interconnected and complex. Various domains like automotive, production, but also critical infrastructures like electronic government systems naturally have to provide communication interfaces to provide more smart and usable functionality. It is therefore crucial to create methodologies to verify the correctness and security of these interfaces. Since those systems are mostly proprietary, verifying their security has become an increasingly laborious task that is afflicted with a high degree of uncertainty. Many approaches rely on static, pre-defined test suites that do not take any architectural or behavioral information into account. Therefore, validation of industrial systems and critical infrastructure often include the use of formal methods that provide more rigor in verification [1]. Such methods, including model checking, often come with the disadvantage of a high effort (especially with modeling systems) in conjunction with being hard to apply at black box systems. To mend these disadvantages, we try to automate the model creation using active automata learning. This technique allows for an automated model generation to be used in a checking process, and is also a technique that is destined for black box systems [2]. We therefore combine the static model checking with dynamic active learning. Furthermore, we want to be able to not only verify the systems, but also correct errors and simulate different behavior. We can also use this technique to create corner-cases for verification by altering the system behavior to a Maximum Credible Accident (MCA) scenario [3]. We therefore follow the approach to translate the learned models into a modeling language (see Sections 2.4 and 5.2).
1.1 Contribution
This paper combines formal methods, namely active automata learning of existing systems and model checking, to create a tool chain for checking cyber-physical systems (particularly communication protocols) for their correctness and security. Our main contributions are:
  • 1.
    An approach for annotating learned models (Mealy machines [4]) with propositions using context-based proposition maps (CPMs - see Section 5.1), effectively creating Kripke structures [5] (combined into a single model with the Mealy machine as a hybrid annotated Mealy machine),
  • 2.
    A formally defined and implemented template that allows for creating Rebeca models [6] from annotated Mealy machines,
  • 3.
    Generalized properties to check security requirements along with a easily usable, scalable method for concretization using CPMs,
  • 4.
    Two case studies of communication protocols (NFC [7] and UDS [8]) demonstrating the applicability and versatility of the approach.
1.2 Approach
We utilize active automata learning via the LearnLib Java library to infer Mealy machine models of a respective System-under-learning (SUL). We therefore created dedicated protocol adapters and defined input alphabets (Section 4). Since this model does not intrinsically contain any properties, we annotate it with sensible propositions. For that, we define using context-based proposition maps (CPMs) that define when propositions become true or false (Section 5.1). We created a tool that uses these maps to annotate Mealy machines represented in the GraphViz format, receiving an combined annotated Mealy machine. We then convert this structure into Rebeca code and verify it to be faithful to the original model (Sections 5.2 and 5.6). Once this step has been performed, we utilize the Rebeca Model Checking tool, RMC [9] to check the models for properties defined in linear temporal logic (LTL). These properties are partly specific to the scrutinized protocol, but can also partly be generalized to check very diverse protocols for the same properties. Eventually, if a property violation is found, the trace can be concretized via the learner’s interface and be used as a test case for verification on the live system. If this confirms the violation, that indicates an issue. Otherwise, the different behavior serves as a counterexample for the learner in a new iteration of the process. That way, we implemented all the phases for going through learning, checking, and testing real-world examples. We automated the complete process by combining (a) the learner(s) using LearnLib [10] (Section 4), (b) a tool for generating Kripke structures and translating it into a Rebeca model to be checked (Section 5) and (c) the Rebeca model checking tool [9], used with generic security properties (Section 6).
The remainder of this paper outlines as follows: Section 2 contains basics, while Section 3 outlines general security requirements to be checked. Section 4 outlines the learner interfaces, Section 5 the Rebeca code generation including a formally defined template and a formal description of the CPM concept. Section 6 the model checking, including property definition. Section 7 briefly describes the relation to testing. Section 8 presents the case studies, Section 9 the related work and Section 10 concludes the paper.
2 Background
In this section, we briefly describe the basics of automata, automata learning, model checking, as well as the frameworks we used and protocols used in the case studies.
2.1 Automata and Kripke Structures
Automata, or state machines, are a fundamental concept of computer science  [11, 12]. To describe cyber-physical systems, we use a certain type of automata, namely Mealy machines [4]. A Mealy machine $\mathcal{M}$ is defined as $\mathcal{M}=(Q,\Sigma,\Omega,\delta,\lambda,q_{0})$, with $Q$ being the set of states, $\Sigma$ the input alphabet, $\Omega$ the output alphabet (that may or may not be identical to the input alphabet), $\delta$ the transition function ($\delta:Q\times\Sigma\rightarrow Q$), $\lambda$ the output function ($\lambda:Q\times\Sigma\rightarrow\Omega$), and $q_{0}$ the initial state. A state machine can describe a system in a way that for any defined set of inputs (input word), it delivers the correct set of outputs (output word) [13]. While Mealy machines are widely used to model system behavior, Kripke structures [5] are basic structures used for model checking [14]. A Kripke structure is similar to a Mealy machine, as both can be seen as types of Labeled Transition Systems (LTS) [15, 16]. However, they differ in that instead of labeling the transitions, Kripke structures have their states labeled with propositions. Coming from propositional logic, these are logical attributes that can be true or false. Any state in which this attribute is true gains it as a proposition. Each state can have an element of the power set of all propositions. The formal definition of a Kripke structure ($\mathcal{K}$) is $\mathcal{K}=(S,S_{0},R,\mathcal{L})$ with $S$ being the set of states, $S_{0}\subseteq S$ the set of input states, $R\subseteq S\times S$ a transition relation and $\mathcal{L}=S\rightarrow 2^{AP}$ the labeling function that attributes an element of the power set of atomic propositions (AP{AP}) to the states.
2.2 Automata Learning
Automata learning was developed as a method to learn regular languages [17]. A learner (i.e., an algorithm) is allowed to ask two kinds of questions: membership queries and equivalence queries. We then assume a teacher that knows all of the rules that determine well-defined words of that language. Membership queries give examples of input words($\sigma\in\Sigma|W_{\sigma}=\langle\sigma_{1},\sigma_{2}..\sigma_{n}\rangle$), where the teacher answers wether the example is part of the language. After a sufficient amount of queries, the learner infers a hypothesis model of an automaton describing the language (in this case a Deterministic Finite Acceptor – DFA). It may then submit an equivalence query, where it presents the hypothesis to the teacher. The teacher either answers that the hypothesis is correct or presents a counterexample. To fit to real-world cyber-physical systems, approach has been adapted to learn Mealy machines as well, alongside with more efficient algorithms [18]. In this case, membership queries yield corresponding output words ($\omega\in\Omega|W_{\omega}=\langle\omega_{1},\omega_{2}..\omega_{n}\rangle$). One popular field to use automata learning is to learn and test black box systems. In this case there is no teacher available that knows the interna of the examined system (SUL). Therefore, membership queries are just passed on to the live system [^1] We are talking about active automata learning here. It is also possible to take the answers from recorded data in passive automata learning.. Equivalence queries are implemented by a sufficient amount of conformance tests [19]. If all are passed, the hypothesis is accepted of correct. Otherwise, a failed conformance test automatically poses a counterexample. In this work, we use Learnlib [10], a well proliferated Java library for automata learning. It provides a classes for interfacing with SULs, implementations of well-known algorithms (L*, Rivest-Schapire, AAAR, ADT, KV, DHC and TTT, as well as an addon for L#  [20]), and conformance testing strategies (complete depth-bounded exploration, random words, random walk, W-method, Wp-method). Other libaries are AALpy [21] and Libalf [22]).
2.2.1 Interface and Abstraction Layer
To interact with real-world systems, we not only need the input alphabet ($\Sigma$), but also a means to communicate with the SUL. To keep the as universal as possible, we therefore define an intermediate layer that serves two purposes: provide an interface to the SUL (including necessary error correction, timeout-handling, etc.) and abstracting the input layer. The latter, called mapper, is necessary to reduce the input space to a size feasible to be exhaustively explored. The mapper is formally defined as a state machine $A=(\Sigma,\Omega,Q,q_{0},\delta,X,Y,\nabla)$, with $\Sigma$ the abstract input alphabet, $\Omega$ the abstract output alphabet, $Q$ being the set of states, $q_{0}$ the initial state, $\delta$ the transition function ($\delta:Q\times\Sigma\rightarrow Q$), $\lambda$ the output function ($\lambda:Q\times\Sigma\rightarrow\Omega$), $X$ the concrete input alphabet, $Y$ the concrete output alphabet, and $\nabla$ the abstraction function ($\nabla:\Sigma\cup\Omega\rightarrow X\cup Y$. Inversely, the same mapper can convert the received concrete output from the SUL back to abstract output to present the learner by replacing the abstraction function with its inverse $\nabla^{-1}:X\cup Y\rightarrow\Sigma\cup\Omega$. The latter must not necessarily be fully defined in all cases. For many protocols, the learner is able to handle concrete outputs and to learn $\Omega$ on-the-fly. For these cases a partial mapper with $A=(\Sigma,Q,q_{0},\delta,X,\nabla_{in})$ is sufficient where $\nabla_{in}:\Sigma\rightarrow X$. However, for some concrete outputs ($Y\prime\subseteq Y$) the inverse mapper must still be defined ($\nabla_{out}^{-1}:Y\prime\rightarrow\Omega$), since some inputs provoke non-deterministic values in the sense that they would not yield the same output even with the exact equal input sequence, which a non-deterministic Mealy machine learner cannot handle. Examples of such outputs are timestamps, counters, and random numbers (such as nonces).
2.3 Linear Temporal Logic and Model Checking
Linear Temporal Logic (LTL) is an extension of propositional logic [23] that allows for expressing logical temporal [^2] In this context, temporal is not to be confused with timed. LTL only allows statements about the modality and succession order of events to occur, not about a duration of any kind. modalities [24]. That is, LTL combines atomic propositions (attributes of a state that can be true or false) formulas using logical operators and extends them with the following modalities:
  • 1.
    always ($\square$, or $G$ for Globally): the proposition must hold in any subsequent state
  • 2.
    eventually ($\lozenge$, or $F$ for Finally): the proposition must hold in some (i.e., any arbitrary) subsequent state (may or may not hold before)
  • 3.
[…]
  • 4.
    until ($\mathcal{U}$ or $U$): the proposition $A_{1}$ must hold until another defined proposition $A_{2}$ occurs ($A_{1}\mathcal{U}A_{2}$). $A_{1}$ may or may not hold after $A_{2}$ has occurred. In any case, $A_{2}$ has to occur at some point.
A model checker runs through all states of a model and checks if their properties hold [24]. The properties are often expressed in LTL or similar logics like Computation tree logic (CTL) [25] or branching-time logic (CTL*) [26]. In automata-theoretic model checking, a model checker ordinarily creates Büchi automata of both the scrutinized model and the negation of the LTL property to check, and examines if their cross product’s accepted language is empty [19].
2.4 Rebeca
Rebeca is a modeling language with model checking support specifically designed to model and model check cyber-physical (i.e., reactive) systems [6]. Rebeca resembles the Java programing language to ease its usage for engineers. It conceptually uses rebecs to model reactive actors. Rebecs are similar to Java objects, but possess interfaces (message servers) to interact with each other and contain a message queue to process messages from other Rebecs in a FIFO manner. Its IDE, Afra [9], contains a model checker, Modere [27]. Each Rebec can have state variables that can constitute checkable properties. The properties are defined as LTL formula in a separate property file. Apart from the usability for software engineering and modeling and model checking cyber-physical systems, it is possible to convert code into Lingua Franca [28]. Lingua Franca is a coordination language, providing a deterministic concurrency model, that supports C, C++, Python, TypeScript, and Rust as targets [29]. Eventually, a Rebeca code can produce a verified program in one of these languages via Lingua Franca. Formally, a Rebeca model $\mathcal{M}$ is defined as $\mathcal{M}=||_{i\in\mathcal{I}}\text{ }r_{i}$, which is the set of concurrent actors (also called rebec for reactive object) $r_{i}$, with $i\in\mathcal{I}$ being an identifier from the set of all identifiers in the model [30]. An actor is defined as $r_{i}:=\langle V_{i},M_{i},K_{i}\rangle$, where $i$ is a unique identifier, $Vi$ the set of state variables, $M_{i}$ the set of method identifiers (local methods and remotely callable methods called message servers), and $K_{i}$ the list of other actors known to actor $i$. Remote method (i.e., message server) calls between actors occur by sending messages ${msg}=\langle{sendid},i,{mtdid}\rangle$, with ${sendid}\in\mathcal{I}$ being the sender actor’s identifier, $i\in\mathcal{I}$ the receiver actor’s identifier, and $mtdid\in M_{i}$ the identifier of the method called on the sender side.
2.5 Near Field Communication and Integrated Circuit Access
Near Field Communication (NFC) is a communication standard [31, 7] that defines wireless communication for small, powerless embedded systems with low data rates. It defines messages for data transmission (information or I blocks), signaling (supervisory or S blocks), and acknowledgments (receive-ready or R blocks) and the modalities operate communications. The main applications are RFID-tags and wireless chip card access. For the latter, it uses the structures of ISO/IEC 7816-4 [32]. This standard logically segments circuits into applications (comparable to directories in a file system) that can be accessed via Dedicated Files (DFs) and contains Elementary Files (EFs) as data stores. The standard defines commands to access (SELECT) data and applications, as well as manipulate data (READ, WRITE, UPDATE, etc.). It also defines GETCHALLENGE and AUthenTICATE commands to implement challenge-response-based authentication mechanisms that protect sensitive data. Usually, an authentication procedure yields a session key, that is subsequently used to encrypt the access to the protected parts of the system. Every command induces a response that contains an unencrypted two-bytes status code (even if the data itself is encrypted). This code is either 9000 for positive results or 6XXX for various error codes.
2.6 Electronically Machine-Readable Travel Documents
Electronically Machine-Readable Travel Documents (eMRTDs) are a logical data structure to store travel document (e.g., passport) data on chips. The International Civil Aviation Organization (ICAO) standardizes this format in their Doc 9303 part 10 standard [33]. It defines two DF areas LDS1 (containing the eMRTD application) and LDS2 (containing travel records, visa records, and additional biometrics). The eMRTD contains the the Common (CM), the Country Verifying Certification Authorities (CVCA), and the Document Security Object (SOD) EFs, as well as EFs for 16 data groups that contain various types of data, like personal data, document data, biometrics (fingerprint, iris). The latter are more sensitive and therefore additionally protected, while the rest is protected by either the older Basic Access Control (BAC) or the newer PACE (Password Authenticated Connection Establishment). BAC generates a key based on some cryptographic operationes with the passport number, expiration date, and the owner’s date of birth. PACE uses a password to encrypt a nonce, which then is the base for a Diffie-Hellman-Merkle key exchange that creates a session key. Besides there are some conditional files outside of both LDS, namely Attributes/Info (ATTR/INFO), Directory (DIR), Card Access (CA), and Card Security (CS). Figure 1 shows an overview of the ICAO eMRTD schema.
Figure 1: Logical Data Structure of Machine Readable Travel Documents from  [34]. Amber is the master file (MF), Cyan are dedicated files (DF), Blue are Elementary Files (EF), and Green are key files. Solid frames means mandatory files, dashed ones optional files. Solid boxes donate the LDS contexts, dashed black boxes requirements, and dashed red boxes necessary authentication.
2.7 Unified Diagnostic Servcies
Unified Diagnostic Services (UDS) is an ISO standard that specifies services for diagnostic testers to control diagnostic (and further) functions on an on-vehicle Electronic Control Unit (ECU) [8]. Therefore, it defines a variety of services, whereby the concrete elaboration mostly lies in the hands of an implementer. The protocol operates on layer 7 of the OSI reference model. Therefore, it relies on an underlying communication protocol. Although it can use multiple protocols, the most proliferated is the Controller Area Network (CAN) protocol. As a client-server-based protocol it defines requests and responses. The requests contain a service ID, optionally followed by sub-service IDs. Particularly, it defines diagnostic and management services: Diagnostic Session Control (0x10), ECUReset (0x11), SecurityAccess (0x27), CommunicationControl (0x28), TesterPresent (0x3E): a keepalive mechanism,, AccessTimingParameter (0x83), SecuredDataTransmission(0x84), ControlDTCSetting (0x85), ResponseOnEvent (0x86), ResponseControl(0x87). It further defines data transmission services: ReadDataByIdentifier (0x22) , ReadMemoryByAddress (0x23) , ReadScalingDataByIdentifier (0x24) , ReadDataByPeriodicIdentifier (0x2A) , DynamicallyDefineDataIdentifier (0x2C) , WriteDataByIdentifier (0x2E) , WriteMemoryByAddress (0x3D), ClearDiagnosticInformation (0x14), ReadDTCInformation (0x19). Lastly also I/O, routine control, and file transfer functions: InputOutputControlByIdentifier (0x2F) , RoutineControl (0x31), RequestDownload (0x34), RequestUpload (0x35), TransferData (0x36), RequestTransferExit (0x37), RequestFileTransfer (0x38). We will use a set of these services as an input alphabet for learning the UDS state machine (see Section 4.2).
3 Generic Security Requirements
In this section, we take some long established security goals [35, 36, 37] that are also aligned with our case studies’ standards documents (ICAO [38] and UDS [39]) and formulate them as checkable requirements in given-when-then [^3] Using given, when, and then as keywords in italics. format [40]. We use the terms MUST, MUST NOT, SHOULD, SHOULD NOT and MAY as defined by the Internet Engineering Task Force (IETF) [41, 42] [^4] Using the same CAPITALIZED style.. These requirements use attributes, we later (in Section 6.1) use as propositions [^5] Using small capitals style. for defining logic properties.
3.1 Authentication
To establish authenticity as a goal, we must require authentication of a user before accessing protected resources (not every resource must be protected, some may be openly accessible – this depends on the system’s security design). As a requirement , we define it as follows:
Requirement R1:Given the system is not in an authenticated state, when an access operation on a protected resource occurs, then the operation MUST NOT return a positive response.
3.2 Confidentiality
To acquire confidentiality (also called secrecy in cryptographic contexts), most protocols allow for encryption (and may require it under certain conditions). We therefore recommend encryption for protected resource access:
Requirement R2:Given the system is in an arbitrary state, when an unsecured access operation on a protected resource occurs, then it SHOULD NOT be successful.
3.3 Privilege Levels
It is common practice for more complex systems and protocols to define more than one privilege level according to different roles and according privileges of an actor in order to provide access to more critical resources. This also applies for UDS, where this is implemented by different session and security levels [8]. Likewise, for eMRTDs, sensitve data (particularly biometric data like fingerprints and iris scans [43]) have to be additionally protected via terminal authentication according to European extended access control [33, 44].
Requirement R3:Given different privilege levels, when the level of privilege is not sufficient, then a performed operation MUST NOT be successful.
3.4 Key Validity
Another important attribute of secure connections is key validity and recency [^6] To avoid misunderstandings, for this work we use the term to refer to using the most recent key, i.e., no newer key must exist. Since we do not consider a time component in the automata, we do not use recency in the sense of freshness.. An invalid key in this sense is either a wrong key or an old key (not recent [^7] To avoid replay and similar attacks [45, 46]). Obviously, this is a fundament for the prior requirements, since an authentication or encryption that accepts invalid keys is intrinsically flawed.
Requirement R4:Given an authentication OR secure access operation, when an actor is asked for a key, then providing an invalid key MUST NOT be successful.
4 Building the Mealy Machines using Automata Learning
We use Learnlib (see Section 2.2) to create a setup to learn models of different protocol SULs, particularly NFC-based eMRTDs and UDS communications of automotive ECUs. Learnlib provides a variety of learning algorithms of which we use the TTT algorightm [47] for its comparable high performance [18]. We rely on previous work for the learning parts of eMRTDs [34] and UDS [48], respectively. In this work, we annotated the learned models’ states with labels for propositions and translate them into code to allow for model checking. In this section we provide some details of the learning processes for the two protocols. While the learner is the same, the SUL adapters differ due to different input alphabets and different means to access the respective SUL.
4.1 NFC Interface
We use NFC to access and learn models of eMRTDs. This section explains the interface and inputs used to learn models of eMRTDs.
4.1.1 Input Alphabet
The used input symbols for learning eMRTDs contain effectively selecting the dedicated file for LDS1, selecting elementary files (particularly, CA/CVCA, CM, CS/SOD, ATR, DIR, and data groups DG1-16888Selecting CA and CVACA, as well as CS and SOD, respectively, works with the same inputs. The difference is whether the DF for LDS1 is selected or not (see Figure 1)., read binary and update binary [^9] We also tried write binary, search binary, erase binary, and read record. These were not supported on used SULs (error code 0x6D00).. These symbols are used in a plain and a secure (encrypted with a key obtained in an authentication process) version. Furthermore, we defined additionally modified versions of the secured commands both using a wrong key (actually an all-zero key) and an old key from a previous authentication process [^10] Secret keys for secured commands are obtained by successfully carrying out authentication. If no authentication happened in an input trace before using a secured command, an all-zero key is used. The same applies if an old key should be used but none or only a recent one is available.. For authentication, also a basic authentication (BAC) process is triggered via a dedicated input symbol. Since we lack an implementation of another authentication method (e.g., PACE or Terminal Authentication), BAC is the only authentication input we use.
4.1.2 Interface Device
We use a Proxmark3 NFC interface devices to access eMRTDs [49]. This devices is capable of handling communication with a broad variety of different NFC dialects, as well as reading from, writing on, flashing and emulating them. Besides, it allows for crafting arbitrary frames, altering any part of the data stream to specific needs.
4.1.3 SUL Adapter
The Java learner communicates with the SUL via a distinctive class that handles input, reset, etc. This SUL class then uses a socket to transmit the inputs to a C++ based adapter program that translates the input symbols into NFC frames and pushes it to the interface device. We altered both the standard proxmark interface program and firmware for efficiency and stability reasons. As a result, we send and receive complete input and output words instead of single symbols [50].
4.2 UDS Interface
4.2.1 Input Alphabet
We use the following inputs from the UDS standard (with the respective service IDs and sub IDs in parentheses): DefaultSession (10_01), ProgrammingSession (10_02), ExtendedDiagnosticSession (10_03), SafetySystemDiagnosticSession (10_04), Clear Diagnostic Information (PowerTrainDTC - 14), SecurityAccess (27), SecurityAccessWithKey (as response to a seed with a legit key - 27_0A), SecurityAccessWithWrongKey (same as SecurityAccessWithKey, but with a random key), Communication Control (28), Authentication (29), TesterPresent (3E), Secured Data (84), Routine Test (31), ReadDataByIdentifyer (22 with IDs 0xF100, 0xF150, and 0xF180), RequestDownload (34), RequestUpload (35), TransferData (36 - using dummy data), TransferExit (37).
4.2.2 Interface Device
In contrast to NFC, the hardware adapter does not need to be modified. To access the SUL, we use an off-the-shelf Peak Systems PCAN-FD device [51], that provides a standard CAN-FD interface for Linux (where it is registered under e.g., /dev/can0) or Windows systems. We access this interface via a Python-based interface described in Section 4.2.3 to send and receive UDS messages over CAN.
4.2.3 SUL Adapter
Analogous to the NFC interface, the learner uses a distinctive SUL class. This class communicates with the device adapter program via a socket interface and passes the input symbols forward. It calls a respective input handler function in a python script. This handler function creates a raw frame (i.e., the raw UDS hex sequence encapsulated in a CAN message) corresponding to the input symbol. It then sends it over the system CAN interface and returns the respective response frame to the learner. This python interface builds on previous work [48] and was expanded with some error handling procedures, CAN-FD support and a variety of additional input symbols.
5 Generating the Rebeca Code
Once the learner provides a model of the SUL (stored in GraphViz format), it can be transformed to code in Rebeca for model checking. This happens in two phases: annotating the model with propositions and subsequently translating into code. The model checker also needs properties to check in the form of linear temporal logic (LTL). This section describes how to obtain these three components.
5.1 Annotating the Mealy Machine with Propositions using CPMs
Automata learning provides us with a SUL’s Mealy machine (see Section 2.2). In a Mealy machine we only have inputs and outputs to distinguish different states and each state does not include any other information. Our goal is to determine the security of a system based on certain properties and it is more intuitive to specify the properties based on the information about the states. We therefore annotate the states with relevant propositions. These propositions should provide the necessary context for sensible security checking.
We define intuitive and usable propositions for our context for each state (e.g., authenticated) and provide a method to annotate the model with these propositions in a correct way. Since the only way of determining the propositions is through the system’s behavior (i.e., input and output), a straightforward approach is to define a context-based proposition map (CPM) that describes which input and output would make a propositions true or false. Therefore, we define gain (for becoming true) and lose (for becoming false) conditions. These condition sets ($C$) define when a state gains ($C_{g}$) and losses ($C_{l}$), respectively, a proposition and constitute the CPM. We define a condition ($c\in C$) as a triplet of a set of propositions $S_{P}\in 2^{P}$, an input set $S_{\Sigma}\in 2^{\Sigma}$, and an output set $S_{\Omega}\in 2^{\Omega}$ ($\langle S_{P},S_{\Sigma},S_{\Omega}\rangle$) and $P$ as set of all propositions in the annotated Mealy machine. Which means that a condition could apply to multiple properties ($p\in P$) to be set by one multiple inputs ($\sigma\in\Sigma$) producing one of multiple outputs ($\omega\in\Omega$). The set $C_{g}$ is applied to all transition labels of the model. The target state of each transition gains $S_{P}$ if the transition’s input and output equals $\sigma\in\S_{\Sigma}$ / $\omega\in S_{\Omega}$ of $c$ according to the Mealy machine’s output function $\lambda$. To automatically apply the propagation of propositions, we create an inheritance map for all propositions. That is, we determine for each transition in the system whether it would propagate the proposition in question. This is the case for any transition which input/output does not match $\sigma/\omega$ for a respective proposition $p$ in the lose set $C_{l}$. With this inheritance map, we iteratively determine the cascading of propositions. For each proposition, we start at the states that possess the respective proposition according to $Cg$ and annotate all target states of transitions in the propagation map for the proposition with $p$. With the updated model, we repeat that step until no new states are to annotate. Eventually, we have a state machine annotated with all propositions according to the defined conditions. As a result, we have labeled states like in a Kripke structure. Since we still maintain the Mealy machine labels of transitions, this resulting annotated Mealy machine can be seen as combined Mealy-Kripke structure. We therefore formally define the annotated Mealy Machine $\mathcal{MK}$ as $\mathcal{MK}=(Q,q_{0},\Sigma,\allowbreak\Omega,\allowbreak,\delta,\lambda,\mathcal{L})$, with $Q$ being the set of states, $q_{0}$ the initial state, $\Sigma$ the input alphabet, $\Omega$ the output alphabet, $\delta$ the transition function and $\lambda$ the output function of the Mealy machine. Finally, $\mathcal{L}$ denotes the labeling function that attributes propositions $p\in P$ to the states. $\mathcal{L}$ corresponds to the used CPM.
Figure 2: We need temporary propositions to specify certain properties where the properties are related to labels on the transitions and we do not have a corresponding proposition in the state for them. Here is a general example for a temporary proposition, stemming from a transition output. Left is the original state ($q$), already labeled with a proposition (AUTHAUTH). On the right side, we introduce an internal state ($\tau$) to gain a checkable proposition for a successful read (READOKREADOK). It has this temporary proposition and inherits all other propositions from $q$.
[…]
5.2 Mapping the Annotated Mealy Machine into Rebeca Code
After annotating the Mealy machine with propositions, a fully automated procedure implemented in Java translates the resulting annotated Mealy machine to Rebeca code (see a formal definition in Section 5.3). In the annotated Mealy machine we see input and outputs of each transition. In the corresponding Rebeca code we have two actors: environment (an external entity giving input and collecting output) and a system (the entity whose behavior is defined by the Annotated Mealy machine). The environment actor contains a request message server and message servers corresponding to each output in the annotated Mealy machine. In the request message server we non-deterministically send messages corresponding to each possible input symbol to the system. For these messages related to each input symbol there is message server in the system actor. Each of these message servers calls back the message server of the environment actor that corresponds to the respective output in the respective state (by call back we mean sending a return message). We keep track of each state using a state variable in the system. Additionally, the system entertains state variables for the propositions. Every time a proposition changes from one state to another (gaining or losing) it is set to true or false, respectively, along with changing the state and ? calling the appropriate output message server on the environment actor. Additionally, the algorithm considers the temporary propositions. If $C_{\tau}$ defines a temporary proposition, a corresponding state variable is set on the environment actor. However, all temporary propositions are reset (i.e., set to false) in the environment actor’s request message server, which is subsequently called by the output message server. This assures that the temporary proposition is present in a state for the model checker, but not maintained further, as it is not part of any state of the Mealy machine. As a summary, the main part of the Mealy machine’s logic is modeled in the system actor, while the environment actor can be seen as an external input giver to the system.
5.3 Formal Definition of a Rebeca Template
We formalize the translation of the annotated Mealy machine by defining a formal Rebeca template by substituting the annotated Mealy components into a Rebeca model definition (see Section 2.4). We define this Rebeca model with two actors: $\mathcal{M}:=r_{sys}||r_{env}$$r_{sys}$ is for a system actor that obtains the annotated Mealy machine’s behavior and $r_{env}$ for an environment actor that serves as an input giver. We define the system as $r_{sys}:=\langle V_{sys},M_{sys},K_{sys}\rangle$. We further define $V_{sys}:=\{q\}\cup AP$. So $r_{sys}$ holds a state variable with an identifier of the current state ($q\in Q$ from the annotated Mealy machine, initially set to $q_{0}$) and $P$, which is the annotated Mealy machine’s set of atomic propositions, for each of which $V_{:}{sys}$ contains a boolean state variable. The only known actor is the environment ($K_{sys}:=\{r_{env}\}$) and the method identifiers are equal the set of input symbols ($M_{sys}:=\Sigma$). The environment ($r_{env}:=\langle V_{env},M_{env},K_{env}\rangle$) contains the temporary propositions from the CPM as state variables ($V_{env}:=C_{\tau})$) and system as other known actor ($K_{env}:=\{r_{sys}\}$). The set of method identifiers consists the output alphabet and a request method that handles the calls of the system and resets the temporary propositions ($M_{env}:=\Omega\cup\{req\}$). In each of the methods of $r_{sys}$ ($i\in M_{sys}$) two things happen: a) we set the curent state identifier ($q$) along with all properties ($p\in AP$) according to the respective transition in the annotated Mealy machine’s transition function that corresponds to the current state and input symbol $\delta(q,\sigma)$ [^11] Note: the Method $m\in M_{sys}$ has a direct, bijective representation in the set of Input Symbols ($\sigma\in\Sigma$): $M_{sys}\leftrightarrow\Sigma$. Additionally having the state variable, we can therefore directly use the transition ($\delta$) and output ($\lambda$) functions of the annotated Mealy machine.. b) we call an output method in $r_{env}$ ($oinM_{env}$) according to the current sate and the annotated Mealy machine’s output function $\lambda(q,\sigma)$, which means sending the message $\langle r_{sys},r_{env},M_{env}(\omega)\rangle$. The environment actor’s ($r_{env}$) request method (reqreq), resets all state variables (which represent the temporary propositions from $C_{\tau}$) to false and randomly picks one input by calling a method from $M_{sys}$ by sending the message $\langle r_{env},r_{sys},M_{sys}(\sigma)\rangle$, with a randomly picked $\sigma\in\Sigma$. Each of the other methods (mapping to a $\omega$ from the annotated Mealy machine) sets a state variable corresponding to a temporary proposition to true, if there is a respective entry in the CPM ($C_{\tau}$) [^12] Depening on the calling methods, which is submitted as a parameter in the message. and then calls the request method (with the message $\langle r_{env},r_{env},{req}\rangle$) to start another interaction cycle. Lastly, we note that the request method is also called by $r_{env}$’constructor, which starts the execution of the Rebeca model ($\mathcal{M}$.
5.4 Illustrative Example
[…]
Figure 3: A simple annotated Mealy machine example (left) with an UML diagram of its Rebeca representation (right), as outlined in Section 5.2. The black arrows in the UML diagram show an example flow an input of $\sigma_{1}$ in state $q_{1}$, the gray ones a flow of an input of $\sigma_{1}$ in state $q_{2}$.
Listing 1: Rebeca code from the illustrative example from Section 5.2.
reactiveclass Environment(3) {
statevars {
boolean omega2set;
knownrebecs {
System system;
Environment(){
self.req();
msgsrv req() {
omega2set=false;
int data = (0);
switch(data) {
case 0: system.sigma1(); break;
msgsrv omega1(){
self.req();
msgsrv omega2(){
omega2set=true;
self.req();
reactiveclass System(3) {
statevars {
int state;
boolean p;
knownrebecs {
Environment environment;
msgsrv sigma1(){
if(state==0) {
state=1;
p=true;
environment.omega1();
} else
if(state==1) {
state=0;
p=false;
environment.omega2();
main {
Environment environment(system):();
System system(environment):();
5.5 Altering the Model
Since in Rebeca a model’s behavior is defined similar to a programming language, altering it is trivial. Furthermore, Rebeca possess an operator for non-determinism (the ?? operator). On the one hand, we can manually add any arbitrary fault that might be interesting to investigate and examine its impact on keeping the properties. On the other hand, this allows for re-introducing some non-deterministic behavior that is abstracted away during the learning process (see Section 2.2.1[52]. By altering the template mentioned in Section 5.3, we just introduce a timeout server on the sender side and alter each input message server on the receiver to trigger a timeout with a certain probability [^14] Realized with a random (??) pick of an integer. If no timeout occurs, the logic continues. Otherwise, we reset the state to initial and call the timeout function on the sender, which sets a temporary proposition for timeout, to allow for checking for it. Since we can set the timeout probability to an arbitrary value, we can play with different scenarios to be checked. We also can convert the resulting Rebeca code back to a Mealy machine, so we possess a model that can contain timeouts and errors, just as the originally learned model would have [^15] If the resulting Rebeca code contains non-determinism, the resulting model is not technically a Mealy machine, but a Mealy-styled non-deterministic state model, for it defines multiple transitions for a given input per state to represent the different possibilities.. Formally, we alter the Rebeca template by extending $M_{env}$ with a timeout method, that sets a state variable ${timeout}\in V_{env}$ (to perform checks for timeouts if needed) and calls the request method. Further, we alter each method $m_{sys}\in M_{sys}$ to randomly (with a settable parameter determining the probability) reset the state identifier $q$ to the initial state $q_{1}$ and call the timeout method ($\langle r_{sys},r_{env},{timeout}\rangle$), instead of the output method ($M_{env}(\omega)$ derived from $\lambda(q,\sigma)$.
5.6 Verifiying the Code
In order to assure the Rebeca code is faithful to the learned model, we check their equivalence. We generate the full state space of the Rebeca model. For that, we run the model checker without any properties to check, so, the model checker does not stop at a violation. This way the full state space is generated, which we export to a Linear Transition System (LTS) representation in the GraphViz format. Based on a previous work in [53], we then use an algorithm that collapses the state space LTS into a Mealy machine and check the bisimilarity of the converted Rebeca Mealy machine with the originally learned model. The Rebeca model checking tool (RMC) supports Linear Temporal Logic (LTL) to formulate properties. For security checking, these LTL form should describe security attributes.
6 Checking the Model
6.1 Defining Generic Properties
Since the requirements in Section 3 are very generic, we can also define generic properties to check them. The become more specific in conjunction with the CPMs (Section 5.1), which define the meaning of propositions used in the property in the context of a specific protocol [^16] If a proposition used in a property is not defined in the respective CPM, we define it as false($\bot$) for the respective property.. For the sake of readability, we restrict our definitions of access to read operations but write, update, and similar operations can be defined the same way.
6.1.1 Authentication
For Requirement R3.1, we define an authenticated state as a state with the proposition authenticated set to false, a when an access operation on a protected resource occurs, then the operation must not return a positive response (access ok).
Property P1: $\square(\neg{AUTH}\land{PROT}\rightarrow\neg{ACCESSOK})$ [^17] Afra does not support implications in the property file, therefore we use $\square(\neg(\neg{AUTH}\land{PROT})\lor\neg{ACCESSOK})$, instead..
Here, we define AUTH{AUTH} for authenticated, PROT{PROT} for protected, and ACCESSOK{ACCESSOK} for a positive response [^18] Example for non-existing protocols: if we have a protocol with no protected resources, we set $PROT:=\bot$, which means that reading is always allowed according to Property P6.1.1. (i.e., a successful read operation).
6.1.2 Confidentiality
For checking the confidentiality (Requirement R3.2), we assume the presence of unsecured and secured operations. For the access operation as used in Section 6.1.1, we therefore define two exemplary subsets: an unsecured read $UREAD$ and a secured read $SREAD$. Since $READ$ is a superset, in the respective context maps we then define that $READ$ will always be set along with $UREAD$ or $SREAD$ is set (using the same conditions, i.e., $c_{READ}\leftrightarrow c_{UREAD}\lor c_{SREAD}$). Given requirement R3.2, we define that an $UREAD$ may not succeed when $PROT$ is true. Please note that we fail to apply this property to the UDS case, since there is no encrypted communication on the CAN bus that servers as a medium for the UDS protocol [^19] While there are some proprietary and research solutions for encrypted CAN, none of them is widely adopted (i.e., used by a larger number of major OEMs). An industry standard gaining momentum, the AUTOSAR Secure On-board Communication (SecOC) [54], only provides authentication and integrity, but not confidentiality..
Property P2: $\square({PROT}\rightarrow\neg\text{ }{UREADOK})$
6.1.3 Privilege Levels
We assume that there are areas that go beyond a basic protection level and that there additional protection measures (e.g., different type or level of authentication) have to be taken (Requirement R3.3). We define this higher protection level as critical resources ($CRIT$) and the additional authentication as privileged ($PRIV$). We further assume that since it is a higher security level privileged authentication ($PRIV$) also includes normal authentication ($AUTH$).
Property P3: $\square(({PRIV}\rightarrow{AUTH})\land(\neg{PRIV}\land{CRIT}\rightarrow\neg{ACCESSOK}))$
6.1.4 Key Validity
A property for using a valid key (Requirement R3.4) is straightforward, since it only has to define that an invalid key must not be accepted in the entire system.
Property P4: $\square(\neg{INVKEYOK})$
Depending on the protocol, we use propositions like $WRONGKEYOK$ or $OLDKEYOK$ that would both be seen as invalid keys.
6.2 Checking the Properties
This section defines and explains the CPMs for eMRTDs and UDS, along with defining some protocol-specific properties that allow for checking the protocol flow.
6.2.1 eMRTDs
For eMRTDs, we define $PROT$ as $DF$ input in $C_{g}$ of the CPM for eMRTDs (Table 1), since the ICAO standard defines the applications to be in a protected zone [^20] Note: we only set DF for selecting LDS1, which corresponds to the eMRTD application. Technically, DF would also be set when selecting other applications (i.e., LDS2). However, we then would design the CPM to distinguish between DFs. This is not necessary in our case, since we only use one., while $AUTH$ is gained by performing a basic authentication (BAC) operation. As $ACCESSOK$ we define a successfully performed file selection in $C_{\tau}$. This means that for P6.1.1 a BAC must be performed before selecting any file inside an application.
For $UREADOK$, we define an unsecured read binary operation. This means that for P6.1.2, only secured read binary operations are allowed.
We define $CRIT$ as a successful selection of data groups 2 or 3 in $C_{\tau}$. These contain biometric data, which is why the ICAO standard requires and additional (particularly terminal) authentication. Since, we do not have an additional authentication method implemented, $PRIV$ is set to false ($\bot$). This means that P6.1.3 only holds when these sensitive data groups cannot be selected in the complete model.
Lastly, we define $INVKEYOK$ for P6.1.4 as any successful secured operation that has been carried out using an old key or an all-zero key in $C_{\tau}$. We realize these by dedicated input symbols (see Section 4.1.1).
We also use a couple of complimentary properties for checking eMRTDs. SecureRead: $\square(\neg(SREADOK\land\neg(DF\land AUTH\land EF)))$, which reads as a secure read operation (SREADOK) can only be successful if authenticated and a protected resource is selected ($DF\land EF$). This also creates the need for a normal read operation not to work on protected resources: PlainRead: $\square(\neg(UREADOK\land(\neg EF\lor DF)))$, which reads as a successful plain read (READOK) cannot happen without a selected resource outside of LDS1 ($\neg EF$$\lor DF$). We can further specify a secure select process with the following property: SecureReadFollowsSecureSelect: $((\neg SREADOK)\ \mathcal{U}\$$\text{SSELEFOK})\lor\square(\neg SREADOK)$. Table 1 gives an overview of the CPM.
6.2.2 UDS
Authentication in UDS works over SecurityAccess and subsequent SecurityAccessWithKey. We therfore define the $AUTH$ proposition after the second step has been successfully exectued. As a canary for P6.1.1, we perform a reading operation on a protected resource. We choose to start the Check Programming Dependencies (0xFF01) routine. This routine does not alter the firmware and is standardized [8]. Furthermore, it is industry practice that this route is protected by a Security Access routine [55]. Therefore we set both $PROT$ and $ACCESSOK$ as a $c_{\tau}$ when starting this routine. For P6.1.2, set perform a direct memory read operation to set $UREADOK$. Unfortunately, we don not have the knowledge of which memory are actually protected (as a workaround defining the CheckASWBit routine could also be defined as $PROT$). For P6.1.3, we define the RequestDownload routine as $CRIT$, as it allows to actually flash the device, while $PRIV$ his defined as SecurityAcccess with a higher level. Lastly for P6.1.4, $INVKEYOK$ is defined as SecurityAcccess with a wrong key (as an own input symbol). Table 2 shows the CPM for UDS.
7 Testing
With a property violation discovered by the learning and checking procedure, the only thing left is to verify the finding on the actual system. Due to relying on conformance testing for equivalence queries, the automata learning process maintains a residual risk for an inaccurate model. We therefore test the trace of a found violation with the life system to exclude false positives. If the actual system behavior matches the one predicted by the model, the violation is confirmed. Otherwise, the trace can be fed back to the learning system to refine the model. The verification of found traces is trivial, as the same tool set from the learning can be used for concretizing the input and interfacing with the SUL. Therefore, this combination can be seen as method for test case generation.
Table 1: CPM for eMRTDs.
Proposition Input Output
Gains ($C_{g}$)
AUTH BAC 9000
DF, PROT DF* 9000
EF EF* 9000
CRIT EF_DG2, EF_DG3 9000
PRIVa TA 9000
Losses ($C_{l}$)
EF, AUTH, PRIV, CRIT DF 9000
AUTH, PRIV EF*, *BIN, *REC 6*
CRIT EF* 9000
Temporary ($C_{\tau}$)
UACCESSOK, ACCESSOK SEL_EF* 9000
SSELEFOK, SACCESSOK, ACCESSOK SSEL_EF* 9000
UREADOK, READOK RD_BIN 9000
SREADOK, READOK SRD_BIN 9000
SSELEFOK SSEL_EF* 9000
INVKEYOK, WRONGKEYOK WS* 9000
INVKEYOK, OLDKEYOK OS* 9000
  • a
    Since, we do not have a Terminal Authentication implementation, this condition is hypothetical (will not be triggered).
8 Evaluation
To show the practical usability and versatility of the described process, we evaluate it on two devices (an eMRTD and an automotive control unit), each speaking one of the communication protocols described earlier in this paper (NFC and UDS). Though the use cases are very different, in both cases the model checking is very similar. At the present abstraction level, we check generic security requirements (Section 3) using according LTL properties (Section 6.1).
8.1 Electronically Machine-Readable Travel Document
For the eMRTD use case, we scrutinize two different Austrian passports, on elder, expired and one current (see a process based on equivalence checking in [34]). The model was learned via LearnLib using the TTT algorithm (with binary closure) and a minimum input trace length of 40 symbols, a maximum of 50, and 150 random walk conformance tests as equivalence oracle.
Figure 4: Simplified version of the annotated eMRDT. Self-loops that do not add to the understanding have been removed for readability. (S)SEL_EF* (ex) and (noex) denotes all elementary file selections of existing and non-existing files.
Using the method described above, we translated the Mealy Machine into Rebeca code (see Listings LABEL:lst:NFCsend and LABEL:lst:NFCrec for parts of the model code for the environment and system agent, respectively). Using the CPM and properties as described in Section 6.2.1, we were able to verify that described security properties hold the scrutinized systems.
Listing 2: Example of generated eMRTD environment actor code.
msgsrv req() {
error=false;
ureadok=false;
sreadok=false;
readok=false;
sselefok=false;
selefok=false;
accessok=false;
int data =?(0,1,2,[..],56);
switch(data) {
case 0: system.pp_sel_ef_ca_cvca(); break;
case 1: system.pp_sel_df_lds1(); break; [..]
case 22: system.pp_rd_bin(); break; [..]
case 28: system.pp_bac(); break; [..]
case 32: system.pp_ssel_ef_dg1(); break;[..]
msgsrv req_9000(int data){
switch(data) {
case 0: accessok=true;selefok=true;
uaccessok=true;break;
case 1: break; [..]
case 22: readok=true;ureadok=true;break; [..]
case 28: break; [..]
case 32: accessok=true;sselefok=true;
saccessok=true;break; [..]
self.req();
Listing 3: Example of generated eMRTD system actor code.
msgsrv pp_rd_bin(){
if(state==0) {
state=0; environment.req_6986();
} else
if(state==1) {
state=1; environment.req_6986();
} else
if(state==2) {
auth=false; state=1; environment.req_6986();
} else
if(state==3) {
state=3; environment.req_9000(22);
} else
if(state==4) {
auth=false; state=5; environment.req_6982();
} else
if(state==5) {
state=5; environment.req_6982();
msgsrv pp_sel_df_lds1(){
if(state==0) {
df=true; prot=true; state=1; environment.req_9000(1);
} else
if(state==1) {
state=1; environment.req_9000(1);
} else
if(state==2) {
auth=false; state=1; environment.req_9000(1);
} else
if(state==3) {
df=true; prot=true; ef=false; state=1; environment.req_9000(1);
} else
if(state==4) {
auth=false; ef=false; state=1; environment.req_9000(1);
} else
if(state==5) {
ef=false; state=1; environment.req_9000(1);
msgsrv pp_bac(){
if(state==0) {
state=0; environment.req_6985();
} else
if(state==1) {
auth=true; state=2; environment.req_9000(28);
} else
if(state==2) {
state=2; environment.req_9000(28);
} else
if(state==3) {
state=3; environment.req_6985();
} else
if(state==4) {
state=4; environment.req_9000(28);
} else
if(state==5) {
auth=true; state=4; environment.req_9000(28);
msgsrv pp_ssel_ef_dg1(){
if(state==0) {
state=0; environment.req_6988();
} else
if(state==1) {
state=1; environment.req_6988();
} else
if(state==2) {
ef=true; state=4; environment.req_9000(32);
} else
if(state==3) {
state=3; environment.req_6988();
} else
if(state==4) {
state=4; environment.req_9000(32);
} else
if(state==5) {
state=5; environment.req_6988();
8.2 Automotive Electronic Control Unit
Our example use case for UDS is an automotive electronic control unit (ECU) from a major European Tier-1 supplier running in a vehicle from a Chinese car manufacturer. The model was learned via LearnLib using the TTT algorithm with a minimum input trace length of 20 symbols, a maximum of 50, and 50 random walks as equivalence oracle. Figure 5 shows the annotated Mealy machine (using the CPM in Table 2) of this ECU.
Table 2: CPM for UDS (SA means SecurityAccess).
Proposition Input Output
Gains ($C_{g}$)
AUTH SAWithKey 67
EXT Extended 5003
PROG Programming 5002
PRIVa HLSAWithKey 67
Losses ($C_{l}$)
EXT,PROG Default 5001
EXT Programming 5002
PROG Extended 5003
AUTH SA, SAwKey, SAwWrongKey 7f
AUTH Session 50
Temporary ($C_{\tau}$)
INVKEYOK, WRONGKEYOK SAwWrongKey 67
PROT CheckASWBit 71
ACCESSOK CheckASWBit 71
UACCESSOK CheckASWBit 71
CRIT RequestDownload 74
UREADOK Read* 62
  • a
    Since, we do not have higher level security access, this condition is hypothetical (will not be triggered).
Figure 5: Simplified annotated Mealy machine of the learned ECU model. Self-loops that do not add to the understanding have been removed and the output truncated for readability. The star (*) denotes any other input not explicitly stated.
We subsequently use the properties from Section 6.1 to check the model for security properties, along with the CPM for UDS (Section 6.2.2). We could verify P6.1.1 . For P6.1.2, we lack the knowledge which memory areas are actually protected. However, it is well known that the CAN bus (over which UDS communication runs) is completely unencrypted, while principally critical (safety) data can be exchanged [56]. Therefore, R3.2 is not met by the protocol itself, regardless of any implementation. To test P6.1.3, we lack a higher layer security access. However, scrutinizing critical functions (RequestDownload), we saw P6.1.3 to hold, as it were not possible to trigger with our current security level. P6.1.4 was found to be violated [^21] According to the standard, wrong key should not be acknowledged, even though a wrong key does not mean the security level must be locked (de-authentication) [39]., as in the (authenticated) state S6 (see Figure 5), the system accepts a wrong key [^22] An issue known to us from previous works [48]..
9 Related Work
This work is based on previous work on model learning and model checking [48, 34, 53]. Peled et al. [19] have provided a very influential paper regarding black box testing that combines automata learning with model checking. They learn a model (using Angluin’s algorithm) create a cross product with a Büchi automaton created from the negate of an LTL property to check. If the accepted language is empty, the property is satisfied. Otherwise, the contained (non-empty) words it serve as a counterexamples that will be tested against a true system model. If these tests yield the same results, the counterexample poses a property violation. Otherwise, it serves as a counterexample for refining the model. They propose both an off-line and on-the-fly approach. The former uses a fully learned (i.e., a full run-through of Angluin’s algorithm) model that is checked. The latter incorporates the checking into the learning process and stops it, if a violation is found, potentially greatly reducing the runtime. Groce et al. [57] advance the former approach by integrating the model checking deeper in the learning algorithm for an efficiency boost. Shijubo et al [58] expand Peled’s approach by introducing strengthened specifications, that tend to find counterexamples that earlier detect divergence between the learned and the actual model, boosting the learning process’ efficiency. We refrain from using on-the-fly checking because we want to fully investigate the model for any violation of a defined set of properties, and therefore not stop at the first found violation, as the purpose is to investigate already implemented real-world systems for all incorrect behavior.  [^23] While the Rebeca model checker also stops with a counterexample when a property violation is found, we can, in contrast to on-the-fly-checking, check the other properties as well on the complete model and therefore determine if the system violates multiple properties. Furthermore, many properties do not actually require a modal operator – e.g., Property 6.1.1 uses a single $\square$ operator, which means it should hold globally and a further distinction is not necessary. We can check these with a simple script running over a Rebeca-generated full state space, enabling us to find all occurrences of a violation. While our work is based on a similar idea as the off-line variant, we extended the approach by annotating Mealy machines with atomic propositions creating Kripke-like structures that we turned into Rebeca code for model checking. That way, we implemented all the phases for going through learning, checking, and testing real-world examples. In our case, however, we automated the complete process by combining (a) the learner(s) (using LearnLib [10]), (b) an automated way of generating Kripke structures and translating it into a checkable model (described in Section 5) and (c) a model checker translates this model and (the negation) LTL properties to Büchi automata, creates a product, and check its language for emptiness [9] with generic properties (Section 6). This improves both the usability of the approach and its applicability to real-world (cyber-pyhsical) systems. Neider and Jansen used a symbolic approach to learn and model-check DFAs (but not Mealy or More machines) [59]. Similarily, Fiterau-Brostean et al. [60, 61] also used a symbolic approach (using the NuSMV model checker). Translating the learned model to Rebeca, our approach in contrast provides more advanced possibilities to manipulate the model, including re-introducing abstracted (e.g. non-deterministic) behavior (see Section 5.5), as they also stated in their work that ironing out timing issues were a major engineering problem. Altering the Rebeca template, it is trivial to re-introduce timeouts (non-deterministically) to check a protocol’s behavior under such conditions. This allows for more advanced checking possibilities. Furthermore, our approach using CPMs does not only allow for a more explicit and comprehensible annotation of states with propositions, but also using generic properties and a clean separation between state proposistions (visible in the annonated Mealy machine) and pure transition outputs (in $C_{\tau}$ and $\tau$ states).
Other approaches [62, 15, 63] combine automata learning with bisimulation-based equivalence checking, which need a specification model to compare the learned automaton with. Our approach needs generalized properties. We do not know of an approach that translates a learned model into a modeling language allowing for both direct property checking and manipulating the model to simulate scenarios.
10 Conclusion
In this paper, we presented an approach to combine automata learning with model checking for cyber-physical systems. We defined Context-based Proposition Maps (CPMs) that provide an annotation mechanism to annotate Mealy machines inferred by Automata Learning with propositions. As a result, we received annotated Mealy machines that combine attributes of Mealy machines and Kripke structures. We subsequently translated the annotated Mealy machine into Rebeca code by formally defining a template using two actors: a receiver modeling the learned system and a sender modeling an external actor that interacts with the system. Starting from four high-level security requirements (authentication, confidentiality, privilege levels, and key validity), we defined generic LTL properties to check models for security. Protocol-specific CPMs provided the context to assure the generated Rebeca code provides all propositions that are used in these properties. This way, each property can be checked with regard to the protocol the examined cyber-physical system runs on. We presented a case study with systems from different domains (a passport and and automotive control unit) speaking different protocols (NFC and UDS) to show the versatility of the approach. We were able to verify the security properties for all requirements on to different passports. On the automotive control unit, we could verify two of them (authentication and privilege levels), while one of them (confidentiality) is intrinsically not provided by the protocol and the last one (key validity) was not met (property violated) on the examined unit.
10.1 Outlook
Further research directions to lift the approach to a larger scale lie in creating adapters for other different protocols and creating other generalized LTL properties to check different aspects of security. Additionally, we also investigate utilizing an LLM to generate LTL properties from threat models to check their respective implementations’ Mealy machines. Also we can improve the learning process by using timing [64] and by extending the model with unknown inputs using input symbol mutation; if the unknown input triggers new behavior, we can dynamically extend the input alphabet (as described here [65]). This allows for a more holistic testing.
11 Acknowledgement
This work has received funding from the European Union under the Horizon Europe programme under grant agreement No. 101168438 (project INTACT) as well as from the Swiss State Secretariat for Education, Research and Innovation and UK Research and Innovation under grant No. 10126241. We further acknowledge the support of the Swedish Knowledge Foundation via the industrial doctoral school RELIANT under grant No. 20220130. The authors want to thank the anonymous reviewers of the Computers & Security Journal, as well as of the International Conference on Runtime Verification for their valuable input. We further want to thank Thomas Grandits for helping to formulate requirements and properties more understandable.

### v2

@@
-Learn, Check, Test – Security Testing Using Automata Learning and Model Checking
+arXiv:2509.22215v2[cs.CR] 27 Feb 2026
 Abstract
-Cyber-physical systems are part of industrial systems and critical infrastructure.
-Therefore, they should be examined in a comprehensive manner to verify their correctness and security.
-At the same time, the complexity of such systems demands such examinations to be systematic and, if possible, automated for efficiency and accuracy.
-A method that can be useful in this context is model checking.
-However, this requires a model that faithfully represents the behavior of the examined system.
-Obtaining such a model is not trivial, as many of these systems can be examined only in black box settings due to, e.g., long supply chains or secrecy.
-We therefore utilize active black box learning techniques to infer behavioral models in the form of Mealy machines of such systems and translate them into a form that can be evaluated using a model checker.
-To this end, we will investigate a cyber-physical systems as a black box using its external communication interface.
-We first annotate the model with propositions by mapping context information from the respective protocol to the model using Context-based Proposition Maps (CPMs).
-We gain annotated Mealy machines that resemble Kripke structures.
-We then formally define a template, which we use to transfer the structures in to a format to be processed by a model checker.
-We further define generic security properties based on basic security requirements (authentication, confidentiality, privilege levels, and key validity).
-Due to the used CPMs, we can instantiate these properties with a meaningful context to check a specific protocol, which makes the approach flexible and scalable.
-Furthermore, the gained model can be easily altered to introduce non-deterministic behavior (like timeouts) or faults and examined if the properties still hold under these different conditions.
-Lastly, we demonstrate the versatility of the approach by providing case studies of very different systems (a passport and an automotive control unit), speaking different communication protocols (NFC and UDS), checked with the same tool chain and the same security properties.
+Security verification of communication protocols in industrial and safety-critical systems is challenging because implementations are often proprietary, accessible only as black boxes, and too complex for manual modeling.
+As a result, existing security testing approaches usually depend on incomplete test suites and/or require labor-intensive modeling, limiting coverage, scalability, and trust.
+This paper addresses the problem of systematically verifying protocol security-properties without access to internal system models.
+We propose a flexible and scalable method for formal verification of communication protocols that combines active automata learning with model checking to enable rigorous security analysis of black-box protocol implementations.
+Behavioral models are first inferred from system interactions using automata learning.
+We then propose Context-based Proposition Maps (CPMs) to enrich the learned models with semantic information, yielding annotated Mealy machines that bridge the gap between learned behavior and property verification.
+These machines are automatically transformed into verifiable models in the Rebeca modeling language, enabling model-checking of generic security properties such as authentication, confidentiality, privilege levels, and key validity, while allowing protocol-specific properties to be manually added.
+The CPMs are also used to populate these generic properties with definitive propositions.
+Furthermore, the gained model can be easily altered to introduce non-deterministic behavior (like timeouts or faults) and be examined if the properties still hold under these different conditions.
+As a result, we gain an automated tool chain that spans from model learning to security property checking of specific protocols.
+The key contributions include: (i) a method for augmenting learned protocol models with security-relevant propositions, (ii) a fully automated transformation pipeline from learned models to model-checking artifacts, (iii) reusable, generic security property templates that are instantiated in protocol-specific models, and (iv) empirical validation through case studies demonstrating applicability in different protocols and domains.
+The results show that the approach enables scalable and systematic discovery of security vulnerabilities in black-box systems while reducing modeling effort and improving automation compared with traditional verification workflows.
 keywords:
@@
 1 Introduction
-Cyber-physical systems in all domains of everyday life become more interconnected and complex.
-Various domains like automotive, production, but also critical infrastructures like electronic government systems naturally have to provide communication interfaces to provide more smart and usable functionality.
-It is therefore crucial to create methodologies to verify the correctness and security of these interfaces.
-Since those systems are mostly proprietary, verifying their security has become an increasingly laborious task that is afflicted with a high degree of uncertainty.
+Interconnected electronic systems in all domains of everyday life become more both ubiquitous and complex.
+Various domains like automotive, production, but also critical infrastructures like electronic government systems naturally have to provide communication capabilities to provide more smart and usable functionality.
+It is therefore crucial to create methodologies to verify the correctness and security of these systems.
+Often those systems are proprietary, which means that verifying their security has become an increasingly laborious task that is afflicted with a high degree of uncertainty.
 Many approaches rely on static, pre-defined test suites that do not take any architectural or behavioral information into account.
@@
 Such methods, including model checking, often come with the disadvantage of a high effort (especially with modeling systems) in conjunction with being hard to apply at black box systems.
-To mend these disadvantages, we try to automate the model creation using active automata learning.
-This technique allows for an automated model generation to be used in a checking process, and is also a technique that is destined for black box systems [2].
-We therefore combine the static model checking with dynamic active learning.
-Furthermore, we want to be able to not only verify the systems, but also correct errors and simulate different behavior.
-We can also use this technique to create corner-cases for verification by altering the system behavior to a Maximum Credible Accident (MCA) scenario [3].
-We therefore follow the approach to translate the learned models into a modeling language (see Sections 2.4 and 5.2).
+To mend these disadvantages, we automate the model creation using active automata learning.
+This technique allows for an automated model generation to be used in a checking process, and is also a technique that is suitable for black box systems [55].
+In this paper, we combine model checking with active learning, introducing novel concepts like Annotated Mealy Machines (AMMs) and Context-Based Proposition Maps (CPMs) to provide the missing link between those two components.
+Furthermore, our technique is able to not only verify the systems, but also help in correcting errors and simulate different behaviors.
+We can use the technique to create corner-cases for verification by altering the system behavior to a maximum credible accident scenario [35].
 1.1 Contribution
-This paper combines formal methods, namely active automata learning of existing systems and model checking, to create a tool chain for checking cyber-physical systems (particularly communication protocols) for their correctness and security.
-Our main contributions are:
+This paper combines formal methods, namely active automata learning of existing systems and model checking, to create a tool chain for checking interactive systems (particularly their communication protocols) for correctness and security.
+In general, formally verifying protocols is a complex and labor-intense task.
+Both creating a faithful system model and sensible, meaningful security properties is time-consuming and error-prone.
+We significantly reduce this workload by automating the model generation through making automatically learned models usable for model checking and by providing templates for generic security properties that can easily be instantiated for concrete protocols.
+The core idea for both are context-based proposition maps (CPMs).
+CPMs are simple, tabular rule sets that are used to define when certain propositions become true or false by actions on the transitions from one state to another.
+For the modeling, we use CPMs to annotate Mealy machines [31], obtained via automata learning, with checkable propositions.
+The result are hybrid structures between Mealy machines and Kripke structures we call annotated Mealy machines (AMMs).
+Creating a formally defined template, also provide an automated translation of AMMs into the Rebeca modeling language, where we can run it through a model checker.
+For the property definition, we provide generic properties (particularly authentication, confidentiality, privilege levels, and key validity) which are represented by composition of propositions that we also define through CPMs.
+If the protocol is known, CPMs are easy to create.
+The only task is to denote which actions in a protocol changes a proposition to true or false.
+This action is represented by an input or output as a label on a transition.
+We call this change of value of a proposition as gain or loss of the proposition in a state when the propositions becomes true or false, respectively.
+In summary, our main contributions are:
   • 1.
-    An approach for annotating learned models (Mealy machines [4]) with propositions using context-based proposition maps (CPMs - see Section 5.1), effectively creating Kripke structures [5] (combined into a single model with the Mealy machine as a hybrid annotated Mealy machine),
+    Proposing context-based proposition maps (CPMs) to be used for annotating the generic models and populating the generic properties with protocol-specific propositions,
   • 2.
-    A formally defined and implemented template that allows for creating Rebeca models [6] from annotated Mealy machines,
+    Offering an approach for automatically annotating learned models (Mealy machines) with propositions using CPMs and creating annotated Mealy machines (AMMS) – hybrid structures of Mealy machines and Kripke structures (suitable for model checking),
   • 3.
-    Generalized properties to check security requirements along with a easily usable, scalable method for concretization using CPMs,
+    Defining and implementing a template that allows for fully automatically creating Rebeca models from AMMs,
   • 4.
-    Two case studies of communication protocols (NFC [7] and UDS [8]) demonstrating the applicability and versatility of the approach.
+    Defining generalized properties to check security requirements along with an easily usable and scalable method for deriving concrete properties based on the generalized ones using CPMs,
+  • 5.
+    Demonstrating the applicability and versatility of the approach using two case studies of communication protocols (NFC [17] and UDS [19]).
 1.2 Approach
@@
 We therefore created dedicated protocol adapters and defined input alphabets (Section 3).
-Since this model does not intrinsically contain any properties, we annotate it with sensible propositions.
-For that, we define using context-based proposition maps (CPMs) that define when propositions become true or false (Section 5.1).
-We created a tool that uses these maps to annotate Mealy machines represented in the GraphViz format, receiving an combined annotated Mealy machine.
-We then convert this structure into Rebeca code and verify it to be faithful to the original model (Sections 5.2 and 5.6).
+Since this model does not intrinsically contain any properties, we annotate it with sensible propositions, creating a hybrid automaton of a Mealy machine and a Kripke structure, which we call an annotated Mealy machine (AMM).
+For that, we define context-based proposition maps (CPMs) that determine the ruleset when propositions become true or false (Section 4).
+We created a tool that uses these maps to annotate Mealy machines represented in the GraphViz format, receiving an AMM.
+We then convert this structure into Rebeca code and verify it to be faithful to the original model in an automated way, formally defining a Rebeca template for that purpose (Sections 6 and 6.5) .
 Once this step has been performed, we utilize the Rebeca Model Checking tool, RMC [23] to check the models for properties defined in linear temporal logic (LTL).
-These properties are partly specific to the scrutinized protocol, but can also partly be generalized to check very diverse protocols for the same properties.
+These properties are partly specific to the scrutinized protocol, but can also partly be generalized to check very diverse protocols for the same properties: we define commonly important properties like authentication, confidentiality, privilege levels, and key validity (Section 7.2), while each protocol might still have constraints that are unique to it (like the order of security-related commands).
 Eventually, if a property violation is found, the trace can be concretized via the learner’s interface and be used as a test case for verification on the live system.
@@
 That way, we implemented all the phases for going through learning, checking, and testing real-world examples.
-We automated the complete process by combining (a) the learner(s) using LearnLib [10] (Section 4), (b) a tool for generating Kripke structures and translating it into a Rebeca model to be checked (Section 5) and (c) the Rebeca model checking tool [9], used with generic security properties (Section 6).
-The remainder of this paper outlines as follows: Section 2 contains basics, while Section 3 outlines general security requirements to be checked.
-Section 4 outlines the learner interfaces, Section 5 the Rebeca code generation including a formally defined template and a formal description of the CPM concept.
+We automated the complete process by combining (a) the learner(s) using LearnLib [21] (Section 3), (b) a tool that generates an annotated Mealy machine and translating it into a Rebeca model to be checked (Section 6) and (c) the Rebeca model checking tool [23], used with generic security properties.
+Figure 1 provides an overview of the proposed automated process of model model checking learned Mealy machines for generic security properties.
+Figure 1: Overview of the proposed approach; boxes denote artifacts and ellipses denote processes or tools.
+The items in amber are artifacts and processes defined in this paper, while boxes in blue are artifacts automatically generated by algorithms defined in this paper.
+Cyan are other items, namely: Mealy machines of SuTs, which are obtained via automata learning and the model checking result which is obtained by running RMC with the respective inputs.
+RMC is shown as an ellipse in Cyan.
+Generic security properties and Rebeca templates are one-time work.
+For each specific problem, only CPMs have to be defined individually.
+The remainder of this paper outlines as follows: Section 2 contains basics.
+Section 3 outlines the automata learner details, Section 6 the Rebeca code generation including a formally defined template and a formal description of the CPM concept.
 Section 7 the model checking, including property definition.
@@
 Automata, or state machines, are a fundamental concept of computer science  [52, 57].
-To describe cyber-physical systems, we use a certain type of automata, namely Mealy machines [4].
+To describe reactive systems, we use a certain type of automata, namely Mealy machines [31].
 A Mealy machine $\mathcal{M}$ is defined as $\mathcal{M}=(Q,\Sigma,\Omega,\delta,\lambda,q_{0})$, with $Q$ being the set of states, $\Sigma$ the input alphabet, $\Omega$ the output alphabet (that may or may not be identical to the input alphabet), $\delta$ the transition function ($\delta:Q\times\Sigma\rightarrow Q$), $\lambda$ the output function ($\lambda:Q\times\Sigma\rightarrow\Omega$), and $q_{0}$ the initial state.
@@
 Each state can have an element of the power set of all propositions.
-The formal definition of a Kripke structure ($\mathcal{K}$) is $\mathcal{K}=(S,S_{0},R,\mathcal{L})$ with $S$ being the set of states, $S_{0}\subseteq S$ the set of input states, $R\subseteq S\times S$ a transition relation and $\mathcal{L}=S\rightarrow 2^{AP}$ the labeling function that attributes an element of the power set of atomic propositions (AP{AP}) to the states.
+The formal definition of a Kripke structure ($\mathcal{K}$) is $\mathcal{K}=(S,S_{0},R,\mathcal{L})$ with $S$ being the set of states, $S_{0}\subseteq S$ the set of input states, $R\subseteq S\times S$ a transition relation and $\mathcal{L}=S\rightarrow 2^{P}$ the labeling function that attributes an element of the power set of (atomic) propositions ($p\in P$}) to the states.
 2.2 Automata Learning
-Automata learning was developed as a method to learn regular languages [17].
+Automata learning, more precisely active automata learning, was developed as a method to learn regular languages [2].
 A learner (i.e., an algorithm) is allowed to ask two kinds of questions: membership queries and equivalence queries.
 We then assume a teacher that knows all of the rules that determine well-defined words of that language.
-Membership queries give examples of input words($\sigma\in\Sigma|W_{\sigma}=\langle\sigma_{1},\sigma_{2}..\sigma_{n}\rangle$), where the teacher answers wether the example is part of the language.
+Membership queries give examples of input words($\sigma\in\Sigma|W_{\sigma}=\langle\sigma_{1},\sigma_{2}..\sigma_{n}\rangle$), where the teacher answers whether the example is part of the language.
 After a sufficient amount of queries, the learner infers a hypothesis model of an automaton describing the language (in this case a Deterministic Finite Acceptor – DFA).
@@
 The teacher either answers that the hypothesis is correct or presents a counterexample.
-To fit to real-world cyber-physical systems, approach has been adapted to learn Mealy machines as well, alongside with more efficient algorithms [18].
+To fit to real-world systems, approach has been adapted to learn Mealy machines as well, alongside with more efficient algorithms [54].
 In this case, membership queries yield corresponding output words ($\omega\in\Omega|W_{\omega}=\langle\omega_{1},\omega_{2}..\omega_{n}\rangle$).
@@
 In this case there is no teacher available that knows the interna of the examined system (SUL).
-Therefore, membership queries are just passed on to the live system [^1] We are talking about active automata learning here.
-It is also possible to take the answers from recorded data in passive automata learning..
+Therefore, membership queries are just passed on to the live system.
 Equivalence queries are implemented by a sufficient amount of conformance tests [37].
@@
 In this work, we use Learnlib [21], a well proliferated Java library for automata learning.
-It provides a classes for interfacing with SULs, implementations of well-known algorithms (L*, Rivest-Schapire, AAAR, ADT, KV, DHC and TTT, as well as an addon for L#  [20]), and conformance testing strategies (complete depth-bounded exploration, random words, random walk, W-method, Wp-method).
+It provides classes for interfacing with SULs, implementations of well-known algorithms (L*, Rivest-Schapire, AAAR, ADT, KV, DHC and TTT, as well as an addon for L#  [25]), and conformance testing strategies (complete depth-bounded exploration, random words, random walk, W-method, Wp-method).
 Other libaries are AALpy [33] and Libalf [6]).
@@
 To interact with real-world systems, we not only need the input alphabet ($\Sigma$), but also a means to communicate with the SUL.
-To keep the as universal as possible, we therefore define an intermediate layer that serves two purposes: provide an interface to the SUL (including necessary error correction, timeout-handling, etc.) and abstracting the input layer.
+To keep it as universal as possible, we therefore define an intermediate layer that serves two purposes: provide an interface to the SUL (including necessary error correction, timeout-handling, etc.) and abstracting the input layer.
 The latter, called mapper, is necessary to reduce the input space to a size feasible to be exhaustively explored.
@@
 2.3 Linear Temporal Logic and Model Checking
-Linear Temporal Logic (LTL) is an extension of propositional logic [23] that allows for expressing logical temporal [^2] In this context, temporal is not to be confused with timed.
+Linear Temporal Logic (LTL) is an extension of propositional logic [41] that allows for expressing logical temporal modalities [4].
+In this context, temporal is not to be confused with timed.
 LTL only allows statements about the modality and succession order of events to occur, not about a duration of any kind.
-modalities [24].
 That is, LTL combines atomic propositions (attributes of a state that can be true or false) formulas using logical operators and extends them with the following modalities:
@@
 2.4 Rebeca
-Rebeca is a modeling language with model checking support specifically designed to model and model check cyber-physical (i.e., reactive) systems [6].
+Rebeca is a modeling language with model checking support specifically designed to model and model check reactive (i.e., communicating) systems [49].
 Rebeca resembles the Java programing language to ease its usage for engineers.
 It conceptually uses rebecs to model reactive actors.
-Rebecs are similar to Java objects, but possess interfaces (message servers) to interact with each other and contain a message queue to process messages from other Rebecs in a FIFO manner.
+Rebecs are similar to Java objects,
+but possess interfaces (message servers) to interact with each other and contain a message queue to process messages from other Rebecs in a FIFO manner.
 Its IDE, Afra [23], contains a model checker, Modere [22].
@@
 The properties are defined as LTL formula in a separate property file.
-Apart from the usability for software engineering and modeling and model checking cyber-physical systems, it is possible to convert code into Lingua Franca [28].
+Apart from the usability for software engineering and modeling and model checking reactive systems, it is possible to convert code into Lingua Franca [46].
 Lingua Franca is a coordination language, providing a deterministic concurrency model, that supports C, C++, Python, TypeScript, and Rust as targets [32].
 Eventually, a Rebeca code can produce a verified program in one of these languages via Lingua Franca.
-Formally, a Rebeca model $\mathcal{M}$ is defined as $\mathcal{M}=||_{i\in\mathcal{I}}\text{ }r_{i}$, which is the set of concurrent actors (also called rebec for reactive object) $r_{i}$, with $i\in\mathcal{I}$ being an identifier from the set of all identifiers in the model [30].
+Formally, a Rebeca model $\mathcal{R}$ is defined as $\mathcal{R}=||_{i\in\mathcal{I}}\text{ }r_{i}$, which is the set of concurrent actors (also called rebec for reactive object) $r_{i}$, with $i\in\mathcal{I}$ being an identifier from the set of all identifiers in the model [48].
 An actor is defined as $r_{i}:=\langle V_{i},M_{i},K_{i}\rangle$, where $i$ is a unique identifier, $Vi$ the set of state variables, $M_{i}$ the set of method identifiers (local methods and remotely callable methods called message servers), and $K_{i}$ the list of other actors known to actor $i$.
@@
 2.5 Near Field Communication and Integrated Circuit Access
-Near Field Communication (NFC) is a communication standard [31, 7] that defines wireless communication for small, powerless embedded systems with low data rates.
-It defines messages for data transmission (information or I blocks), signaling (supervisory or S blocks), and acknowledgments (receive-ready or R blocks) and the modalities operate communications.
-The main applications are RFID-tags and wireless chip card access.
-For the latter, it uses the structures of ISO/IEC 7816-4 [32].
-This standard logically segments circuits into applications (comparable to directories in a file system) that can be accessed via Dedicated Files (DFs) and contains Elementary Files (EFs) as data stores.
-The standard defines commands to access (SELECT) data and applications, as well as manipulate data (READ, WRITE, UPDATE, etc.).
 It also defines GETCHALLENGE and AUthenTICATE commands to implement challenge-response-based authentication mechanisms that protect sensitive data.
@@
 The International Civil Aviation Organization (ICAO) standardizes this format in their Doc 9303 part 10 standard [16].
-It defines two DF areas LDS1 (containing the eMRTD application) and LDS2 (containing travel records, visa records, and additional biometrics).
-The eMRTD contains the the Common (CM), the Country Verifying Certification Authorities (CVCA), and the Document Security Object (SOD) EFs, as well as EFs for 16 data groups that contain various types of data, like personal data, document data, biometrics (fingerprint, iris).
+It defines two Dedicated File (DF) areas LDS1 (containing the eMRTD application) and LDS2 (containing travel records, visa records, and additional biometrics).
+The eMRTD contains the Common (CM), the Country Verifying Certification Authorities (CVCA), and the Document Security Object (SOD) Elementary Files (EFs), as well as EFs for 16 data groups that contain various types of data, like personal data, document data, and biometrics (fingerprint, iris).
 The latter are more sensitive and therefore additionally protected, while the rest is protected by either the older Basic Access Control (BAC) or the newer PACE (Password Authenticated Connection Establishment).
-BAC generates a key based on some cryptographic operationes with the passport number, expiration date, and the owner’s date of birth.
+BAC generates a key based on some cryptographic operations with the passport number, expiration date, and the owner’s date of birth.
 PACE uses a password to encrypt a nonce, which then is the base for a Diffie-Hellman-Merkle key exchange that creates a session key.
@@
 Figure 2: Logical Data Structure of Machine Readable Travel Documents from  [29].
+This shows the applications and files on an eMRTD chip.
 Amber is the master file (MF), Cyan are dedicated files (DF), Blue are Elementary Files (EF), and Green are key files.
@@
 The requests contain a service ID, optionally followed by sub-service IDs.
-Particularly, it defines diagnostic and management services: Diagnostic Session Control (0x10), ECUReset (0x11), SecurityAccess (0x27), CommunicationControl (0x28), TesterPresent (0x3E): a keepalive mechanism,, AccessTimingParameter (0x83), SecuredDataTransmission(0x84), ControlDTCSetting (0x85), ResponseOnEvent (0x86), ResponseControl(0x87).
-It further defines data transmission services: ReadDataByIdentifier (0x22) , ReadMemoryByAddress (0x23) , ReadScalingDataByIdentifier (0x24) , ReadDataByPeriodicIdentifier (0x2A) , DynamicallyDefineDataIdentifier (0x2C) , WriteDataByIdentifier (0x2E) , WriteMemoryByAddress (0x3D), ClearDiagnosticInformation (0x14), ReadDTCInformation (0x19).
-Lastly also I/O, routine control, and file transfer functions: InputOutputControlByIdentifier (0x2F) , RoutineControl (0x31), RequestDownload (0x34), RequestUpload (0x35), TransferData (0x36), RequestTransferExit (0x37), RequestFileTransfer (0x38).
-We will use a set of these services as an input alphabet for learning the UDS state machine (see Section 4.2).
-3 Generic Security Requirements
-In this section, we take some long established security goals [35, 36, 37] that are also aligned with our case studies’ standards documents (ICAO [38] and UDS [39]) and formulate them as checkable requirements in given-when-then [^3] Using given, when, and then as keywords in italics.
-format [40].
-We use the terms MUST, MUST NOT, SHOULD, SHOULD NOT and MAY as defined by the Internet Engineering Task Force (IETF) [41, 42] [^4] Using the same CAPITALIZED style..
-These requirements use attributes, we later (in Section 6.1) use as propositions [^5] Using small capitals style.
-for defining logic properties.
-3.1 Authentication
-To establish authenticity as a goal, we must require authentication of a user before accessing protected resources (not every resource must be protected, some may be openly accessible – this depends on the system’s security design).
-As a requirement , we define it as follows:
-Requirement R1:Given the system is not in an authenticated state, when an access operation on a protected resource occurs, then the operation MUST NOT return a positive response.
-3.2 Confidentiality
-To acquire confidentiality (also called secrecy in cryptographic contexts), most protocols allow for encryption (and may require it under certain conditions).
-We therefore recommend encryption for protected resource access:
-Requirement R2:Given the system is in an arbitrary state, when an unsecured access operation on a protected resource occurs, then it SHOULD NOT be successful.
-3.3 Privilege Levels
-It is common practice for more complex systems and protocols to define more than one privilege level according to different roles and according privileges of an actor in order to provide access to more critical resources.
-This also applies for UDS, where this is implemented by different session and security levels [8].
-Likewise, for eMRTDs, sensitve data (particularly biometric data like fingerprints and iris scans [43]) have to be additionally protected via terminal authentication according to European extended access control [33, 44].
-Requirement R3:Given different privilege levels, when the level of privilege is not sufficient, then a performed operation MUST NOT be successful.
-3.4 Key Validity
-Another important attribute of secure connections is key validity and recency [^6] To avoid misunderstandings, for this work we use the term to refer to using the most recent key, i.e., no newer key must exist.
-Since we do not consider a time component in the automata, we do not use recency in the sense of freshness..
-An invalid key in this sense is either a wrong key or an old key (not recent [^7] To avoid replay and similar attacks [45, 46]).
-Obviously, this is a fundament for the prior requirements, since an authentication or encryption that accepts invalid keys is intrinsically flawed.
-Requirement R4:Given an authentication OR secure access operation, when an actor is asked for a key, then providing an invalid key MUST NOT be successful.
-4 Building the Mealy Machines using Automata Learning
-We use Learnlib (see Section 2.2) to create a setup to learn models of different protocol SULs, particularly NFC-based eMRTDs and UDS communications of automotive ECUs.
+Table 1 shows an overview of some of these services and their identifiers.
+Table 1: Some of the most widely used UDS services.
+Service Identifier
+Diagnostic Session Control 0x10
+ECUReset 0x11
+ClearDiagnosticInformation 0x14
+ReadDTCInformation 0x19
+ReadDataByIdentifier 0x22
+ReadMemoryByAddress 0x23
+ReadScalingDataByIdentifier 0x24
+SecurityAccess 0x27
+CommunicationControl 0x28
+ReadDataByPeriodicIdentifier 0x2A
+DynamicallyDefineDataIdentifier 0x2C
+WriteDataByIdentifier 0x2E
+InputOutputControlByIdentifier 0x2F
+RoutineControl 0x31
+RequestDownload 0x34
+RequestUpload 0x35
+TransferData 0x36
+RequestTransferExit 0x37
+RequestFileTransfer 0x38
+WriteMemoryByAddress 0x3D
+TesterPresent 0x3E
+AccessTimingParameter 0x83
+SecuredDataTransmission 0x84
+ControlDTCSetting 0x85
+ResponseOnEvent 0x86
+ResponseControl 0x87
+We will use a set of these services as an input alphabet for learning the UDS state machine.
+3 Building Mealy Machines using Automata Learning
+We use Learnlib to create a setup to learn models of different protocol SULs, particularly NFC-based eMRTDs and UDS communications of automotive ECUs.
 Learnlib provides a variety of learning algorithms of which we use the TTT algorightm [20] for its comparable high performance [54].
@@
 In this work, we annotated the learned models’ states with labels for propositions and translate them into code to allow for model checking.
+As black-box-based automata learning of real-world systems naturally introduces the possibility of errors into the learned model.
+For that reason reason we treat found violations in the model not as issues but test cases only.
+We then use this test case on the real system.
+If the model behaves wrongly as predicted, the violation was confirmed.
+Otherwise, we gained a counterexample to feed the learner for a refined model.
 In this section we provide some details of the learning processes for the two protocols.
-While the learner is the same, the SUL adapters differ due to different input alphabets and different means to access the respective SUL.
+While the learner is the same, the SUL adapters that provide access to the physical system, differ due to different input alphabets and different means to access the respective SUL.
 3.1 NFC Interface
@@
 3.1.1 Input Alphabet
-The used input symbols for learning eMRTDs contain effectively selecting the dedicated file for LDS1, selecting elementary files (particularly, CA/CVCA, CM, CS/SOD, ATR, DIR, and data groups DG1-16888Selecting CA and CVACA, as well as CS and SOD, respectively, works with the same inputs.
-The difference is whether the DF for LDS1 is selected or not (see Figure 1)., read binary and update binary [^9] We also tried write binary, search binary, erase binary, and read record.
-These were not supported on used SULs (error code 0x6D00)..
+The used input symbols for learning eMRTDs contain effectively selecting the dedicated file for LDS1, selecting elementary files (particularly, CA/CVCA, CM, CS/SOD, ATR, DIR, and data groups DG1-16, read binary and update binary.
+Selecting CA and CVACA, as well as CS and SOD, respectively, thereby works with the same inputs.
+The difference is whether the DF for LDS1 is selected or not (see Figure 2).
+We also tried write binary, search binary, erase binary, and read record.
+These were not supported on used SULs (error code 0x6D00).
 These symbols are used in a plain and a secure (encrypted with a key obtained in an authentication process) version.
-Furthermore, we defined additionally modified versions of the secured commands both using a wrong key (actually an all-zero key) and an old key from a previous authentication process [^10] Secret keys for secured commands are obtained by successfully carrying out authentication.
+Furthermore, we defined additionally modified versions of the secured commands both using a wrong key (actually an all-zero key) and an old key from a previous authentication process.
+Secret keys for secured commands are obtained by successfully carrying out authentication.
 If no authentication happened in an input trace before using a secured command, an all-zero key is used.
-The same applies if an old key should be used but none or only a recent one is available..
+The same applies if an old key should be used but none or only a recent one is available.
 For authentication, also a basic authentication (BAC) process is triggered via a dedicated input symbol.
@@
 We use a Proxmark3 NFC interface devices to access eMRTDs [13].
-This devices is capable of handling communication with a broad variety of different NFC dialects, as well as reading from, writing on, flashing and emulating them.
+This device is capable of handling communication with a broad variety of different NFC dialects, as well as reading from, writing on, flashing and emulating them.
 Besides, it allows for crafting arbitrary frames, altering any part of the data stream to specific needs.
@@
 It then sends it over the system CAN interface and returns the respective response frame to the learner.
-This python interface builds on previous work [48] and was expanded with some error handling procedures, CAN-FD support and a variety of additional input symbols.
-5 Generating the Rebeca Code
-Once the learner provides a model of the SUL (stored in GraphViz format), it can be transformed to code in Rebeca for model checking.
-This happens in two phases: annotating the model with propositions and subsequently translating into code.
-The model checker also needs properties to check in the form of linear temporal logic (LTL).
-This section describes how to obtain these three components.
-5.1 Annotating the Mealy Machine with Propositions using CPMs
-Automata learning provides us with a SUL’s Mealy machine (see Section 2.2).
+This Python interface builds on previous work [10] and was expanded with improved error handling, CAN-FD support and around the double amount of input symbols from the UDS protocol.
+Figure 3: A graphical Mealy machine representation of an example for a simple authenticated file read mechanism.
+4 Context-Based Proposition Maps
+Once the learner provides a model of the SUL (stored in GraphViz format), we create Context-Based Proposition Maps (CPMs) to annotate the model with propositions of our interest.
+The generic requirements we defined in this paper are the basis for an engineer to populate the CPM table.
+We defined generic properties (Section 7.2) which constitute of generic propositions like authentication and successful access.
+The engineer makes these generic propositions specific in the CPM table based on the specific protocol.
+Each state in the Mealy machine is augmented with propositions which will also be included in the Rebeca model and will be used for checking the desired properties while model checking.
+4.1 Motivating Example
+To give an illustrative example, we define an Mealy machine that represents a trivial authenticated file access.
+The machine has two states ($Q=\{S0,S1\}$), two input ($\Sigma=\{$CRED,READ}\}), and two output ($\Omega=\{$OK,ERR}\}) symbols.
+Further, we have the initial state $q_{0}=S0$, the transition function
+[…]
+and the output function
+[…]
+Figure 3 shows a graphical representation of this example.
+This machine essentially describes that before reading, a user must give its credentials ($CRED$) to authenticate and that provding credentials twice de-authenticates.
+We will re-use this example Mealy machine to demonstrate the usage of CPMs to annotate Mealy machines with propositions and to produce Rebeca code.
+4.2 Propositions and Mealy Machines
+Automata learning provides us with a SUL’s Mealy machine just like our example.
 In a Mealy machine we only have inputs and outputs to distinguish different states and each state does not include any other information.
 Our goal is to determine the security of a system based on certain properties and it is more intuitive to specify the properties based on the information about the states.
-We therefore annotate the states with relevant propositions.
+We therefore need to annotate the states with relevant propositions.
 These propositions should provide the necessary context for sensible security checking.
 We define intuitive and usable propositions for our context for each state (e.g., authenticated) and provide a method to annotate the model with these propositions in a correct way.
-Since the only way of determining the propositions is through the system’s behavior (i.e., input and output), a straightforward approach is to define a context-based proposition map (CPM) that describes which input and output would make a propositions true or false.
-Therefore, we define gain (for becoming true) and lose (for becoming false) conditions.
-These condition sets ($C$) define when a state gains ($C_{g}$) and losses ($C_{l}$), respectively, a proposition and constitute the CPM.
-We define a condition ($c\in C$) as a triplet of a set of propositions $S_{P}\in 2^{P}$, an input set $S_{\Sigma}\in 2^{\Sigma}$, and an output set $S_{\Omega}\in 2^{\Omega}$ ($\langle S_{P},S_{\Sigma},S_{\Omega}\rangle$) and $P$ as set of all propositions in the annotated Mealy machine.
-Which means that a condition could apply to multiple properties ($p\in P$) to be set by one multiple inputs ($\sigma\in\Sigma$) producing one of multiple outputs ($\omega\in\Omega$).
-The set $C_{g}$ is applied to all transition labels of the model.
-The target state of each transition gains $S_{P}$ if the transition’s input and output equals $\sigma\in\S_{\Sigma}$ / $\omega\in S_{\Omega}$ of $c$ according to the Mealy machine’s output function $\lambda$.
-To automatically apply the propagation of propositions, we create an inheritance map for all propositions.
-That is, we determine for each transition in the system whether it would propagate the proposition in question.
-This is the case for any transition which input/output does not match $\sigma/\omega$ for a respective proposition $p$ in the lose set $C_{l}$.
-With this inheritance map, we iteratively determine the cascading of propositions.
-For each proposition, we start at the states that possess the respective proposition according to $Cg$ and annotate all target states of transitions in the propagation map for the proposition with $p$.
-With the updated model, we repeat that step until no new states are to annotate.
+Since the only way of determining the propositions is through the system’s behavior (i.e., input and output), a straightforward approach is to define a map to show which input and output would make a proposition true or false.
+We call this map a Context-based Proposition Map (CPM).
+Each row in the CPM table defines a condition.
+Table 2 shows a CPM that we use to annotate the example Mealy machine from Section 4.1 with propositions for Property P7.2.1 in Section 7.2.1 (authentication).
+Table 2: CPM for the Example Mealy.
+Proposition Input Output
+Gains ($C_{g}$)
+AUTH CRED OK
+PROT * *
+Losses ($C_{l}$)
+AUTH CRED ERR
+Implicit State Propositions ($C_{\tau}$)
+ACCESSOK READ OK
+a)Refer to caption Refer to caption b)Refer to caption Refer to caption
+Figure 4: We need implicit state propositions to specify certain properties where the properties are related to labels on the transitions and we do not have a corresponding proposition in the state for them.
+Here is a general example for a implicit state proposition, stemming from a transition output.
+The two diagrams in 4a) show a state that possesses the $AUTH$ proposition, while 4b) show one that does not.
+In both cases, the the original state ($q$) is the left.
+On the right side of each a) and b), we introduce an implicit state ($\tau$) to gain a checkable proposition for a successful read (READOKREADOK).
+It has this implicit state proposition and inherits all other propositions from $q$.
+This shows the purpose of such implicit state propositions.
+Given the property that one has to be authenticated to read a resource, the 4a) diagrams show a valid operation, while the 4b) violate that property inside the $\tau$ state.
+Using automata-based model checking, we cannot check for these cases without implicit state propositions, as we need a state where $READOK$ is set.
+4.3 Conditions and Propositions
+We define a condition as a connection between the propositions of our interest and the transitions in a Mealy machine.
+In each state, some propositions become true and some become false.
+We make the connection between the inputs and outputs on each incoming transition with gaining and losing a proposition (becoming true and false, respectively).
+Thereby, a condition is matched to each transition one-by-one (without a history) – if a condition matches a transition, the target state gets the specified property.
+Definition 1 (Condition).
+We define a condition $c\in C$ as a triplet of a set of propositions $S_{p}\in 2^{P}$, an input set $S_{\sigma}\in 2^{\Sigma}$, and an output set $S_{\omega}\in 2^{\Omega}$ ($\langle S_{P},S_{\sigma},S_{\omega}\rangle$); with $P$ as the set of all propositions to be attributed to a Mealy machine, according to the semantics of the examined system protocol and properties of interest.
+Since we are using power sets, a condition could determine multiple propositions ($p\in P$) to be set by one of multiple inputs ($\sigma\in\Sigma$) producing one of multiple outputs ($\omega\in\Omega$).
+We define a match function for a condition by looking at each transition.
+If the input label of a transition matches one element of the input set of the condition, and analogously the output label of the same transition matches one element of the output set, then the match function returns true.
+Definition 2 (Condition Match Function).
+[…]
+The match function gives us the verdict if a transition (uniquely identified by its origin state $q$ and its input $\sigma$) matches a condition $c$.
+Using this match function, we can determine if a target state of such transition gains additional propositions (i.e.
+setting propositions from false to true) stated in a condition $c$.
+We define that, iff a transition (uniquely identified by $q$ and $\sigma$) matches the condition (i.e., the match function for $c$,$q$, and $\sigma$ returns true), we add the propositions in $S_{P}$ of $c$ to state ($q\prime$).
+Therefore, we define a set of proposition gain conditions as
+Definition 3 (Proposition Gain Conditions).
+[…]
+Analogously, we define a set of loss conditions that determine when a target state $q\prime$ of a transition loses a set of propositions $S_{p}$ (propsitions that become flase).
+Definition 4 (Proposition Loss Conditions).
+[…]
+4.4 Implicit State Propositions
+We may not be able to assign all the propositions in our CPM to a state in the Mealy machine using the above annotation algorithm (process).
+So, we have to create new states (which we call implicit states) and assign those propositions to these implicit states.
+For an example, a state $q$ contains an operation (we want to attribute with proposition $p$) as a self-loop transition.
+Before performing the operation, $p$ should be false, afterwards it should become true.
+In such a case it is undecidable if state $q$ has proposition $p$ or not.
+It might, however, still be needed to check if $p$ does not occur in combination with some other proposition (e.g., a read must only occur if authentication is also set – see Property P7.2.1 in Section 7.2).
+To resolve this, we define implicit or implicit ($\tau$) propositions.
+That is, propositions that are true only in a single, implicit state.
+To realize this, we split each transition (in the example above, the self-loop with $p$) into two parts: one transition leading from $q$ to a newly introduced implicit ($\tau$) state and one transition leading from the $\tau$ state to the original target state.
+We then label the transition towards it with the original input ($\sigma\in\Sigma$) and output ($\omega\in\Omega$) and the transition to the original transition’s target state with an empty input and output ($\sigma\leftarrow\varepsilon,\omega\leftarrow\varepsilon$).
+These two transitions are the only ones connecting the $\tau$ state to the graph.
+This mechanism formally makes the resulting construct a Mealy-style subsequential transducer with partially defined transition and output functions [40].
+Every $\tau$ state has only a defined transition and output function for the emptiness symbol $\varepsilon$ as input, so not every state $q\in Q$ has a defined transition for every input $\sigma\in\Sigma$.
+If the split transition matches a implicit condition ($c_{\tau}\in C_{\tau}$ - see Definition 6), we also label the $\tau$ state with the condition’s proposition set ($P(q_{\tau})\leftarrow c[S_{p}]$).
+Figure 4 gives an illustrative example of this mechanism.
+We therefore define a transition split function that creates a new implicit state, splits the original transition in two parts (from the origin to the implicit and from the implicit to the target state), and copies the implicit state propositions from the origin state.
+We further create a set of implicit states $Q_{\tau}$ that do not contain the externally visible states of the classic Mealy machine part.
+Definition 5 (Transition Split Function).
+The transition split function creates a new implicit state $q_{\tau}$, turns a transition $t(q\xrightarrow[\omega]{\sigma}q\prime)$ into two transitions $t\prime(q\xrightarrow[\omega]{\sigma}q_{\tau})$ and $t\prime\prime(q_{\tau}\xrightarrow[\varepsilon]{\varepsilon}q\prime)$, and labels $q_{\tau}$ with the propositions from $q$: $f_{\tau}(q,\sigma,\omega,q\prime):=(Q\leftarrow Q\cup\{q_{\tau}\},Q_{\tau}\leftarrow Q_{\tau}\cup\{q_{\tau}\},T\leftarrow T\setminus\{(q,\sigma,\omega,q\prime)\}\cup\{(q,\sigma,\omega,q_{\tau}),(q_{\tau},\varepsilon,\varepsilon,q)\},P(q_{\tau})\leftarrow P(q)$.
+For the structure with split transitions, we then define a set of implicit state proposition gain conditions that will add a set of extra propositions $S_{p}$ to newly introduced $\tau$ states (we check this by using the implicit state set $Q_{\tau}$, we just introduced above) if there is a matching implicit condition $c_{\tau}$, similar to the standard proposition gain conditions.
+Definition 6 (Implicit State Proposition Gain Conditions).
+We define the set of implicit state proposition gain conditions $C_{\tau}$ as a set of conditions $c$ where the target state of a transition $q\prime=\delta(q)$ gets additional properties $S_{p}$ defined in $c$ (relative to the origin $q$) if the condition matches the transition ($\mathrm{match}_{c}(q,\sigma)=\top$) and that target state is an implicit state ($q\prime\in Q_{\tau}$) $C_{\tau}:=\{c\in C\mid\tau_{c}\}$ and $\tau_{c}(q,\sigma):=\begin{cases}P(\delta(q))\leftarrow P(\delta(q))\cup S_{p}&\text{if }(\mathrm{match}_{c}(q,\sigma)\land\\ &(\delta(q)\in Q_{\tau})),\\ P(\delta(q))\leftarrow P(\delta(q))&\text{otherwise.}\end{cases}$
+That means that the same mechanism as in $C_{g}$ applies to $\tau$ states via $C_{\tau}$.
+Since the implicit state propositions should only be valid in the implicit state explicitly mentioned in $C_{\tau}$, we do not need loss conditions for implicit state propositions.
+4.5 CPM Composition
+Finally, we define a Context-based Proposition Map as a set of the three formerly defined sets.
+Definition 7 (Context-based Proposition Map).
+A Context-based Proposition Map (CPM) is the set of the three sets for proposition gain conditions $C_{g}$, proposition loss conditions $C_{l}$, and implicit state proposition gain conditions $C_{\tau}$ $C:=\{C_{g},C_{l},C_{\tau}\}$
+We use these proposition maps (cf. Figure 1) to a) label Mealy machines with propositions (Section 5) and b) populate generic properties with protocol-specific propositions (Section 7.2).
+5 Annotating Mealy Machines with Propositions
+Since the Mealy machines used in our approach lack propositions that are needed for (Kriepke structure-based) model checking, we use a CPM to annotate them.
+To this end, we first apply the conditions in set $C_{g}$ to all transition labels of the model.
+The target state of each transition gains $S_{P}$ if the transition match $c$ according to matching function (Definition 2).
+We further assume that propositions will propagate over transitions, unless there is a condition that prevents this.
+For instance, a system that is in an authenticated state will stay so, even if the state changes, until there is an event that triggers a de-authentication.
+These preventive conditions are the ones we denote in the proposition loss condition set $C_{l}$.
+We therefore, for each transition $t\in T$, add the propositions of the origin state to the ones of the target state, except if the transition matches an element of the set of lose conditions: ($\forall t\in T\mid P(q\prime_{t}):=P(q\prime_{t})\cup P(q_{t})\text{ if }\neg\text{ loss}_{c}(t[q],t[\sigma])$).
+Since now more states have a certain proposition set, this proposition may propagate even further.
+We therefore reiterate this process with the updated model until no state gains a new proposition, which means the propagation is complete.
+Algorithm 1 shows a pseudocode of the labeling algorithm to annotate Mealy machines with propositions.
+Algorithm 1 Mealy Machine Annotation
+1://Step 1: Parse Mealy Machine
+2:for all lines $\ell$ in DOT file of the Mealy machine do
+3:  if $\ell$ denotes state then
+4:   $Q\leftarrow Q\cup(q\text{ from }\ell)$
+5:   if $\ell$ declares initial state then
+6:     $q_{0}\leftarrow$ ($q$ from $\ell$)
+7:   end if
+8:  end if
+9:  if $\ell$ denotes transition then
+10:   $t\leftarrow$ ($(q,q\prime,\sigma,\omega)$ from $\ell$); $T\leftarrow T\cup t$
+11:  end if
+12:end for
+13://Step 2: Distribute Initial Propositions using the Gain Conditions
+14:for all transitions $t$ in $T$ do
+15:  for all conditions $c_{g}$ in $C_{g}$ do
+16:   if match($c_{g},t$) then
+17:     $(q,i,o,q\prime)\leftarrow t,P(q\prime)\leftarrow P(q\prime)\cup c[S_{p}]$
+18:   end if
+19:  end for
+20:end for
+21://Step 3: Iteratively Propagate Propositions unless there is a Loss Condition
+22:repeat
+23:  changed $\leftarrow\bot$
+24:  for all transitions $t$ in $T$ do
+25:   for all conditions $c_{l}$ in $C_{l}$ do
+26:     for all properties $p$ in $P(t[q])$ do
+27:      if $(\neg(p\subseteq P(q\prime))\land(\neg match(c_{l},t))$ then
+28:        $P(q^{\prime})\leftarrow P(q^{\prime})\cup P(q)$
+29:        changed $\leftarrow\top$
+30:      end if
+31:     end for
+32:   end for
+33:  end for
+34:until not changed
 Eventually, we have a state machine annotated with all propositions according to the defined conditions.
 As a result, we have labeled states like in a Kripke structure.
-Since we still maintain the Mealy machine labels of transitions, this resulting annotated Mealy machine can be seen as combined Mealy-Kripke structure.
-We therefore formally define the annotated Mealy Machine $\mathcal{MK}$ as $\mathcal{MK}=(Q,q_{0},\Sigma,\allowbreak\Omega,\allowbreak,\delta,\lambda,\mathcal{L})$, with $Q$ being the set of states, $q_{0}$ the initial state, $\Sigma$ the input alphabet, $\Omega$ the output alphabet, $\delta$ the transition function and $\lambda$ the output function of the Mealy machine.
-Finally, $\mathcal{L}$ denotes the labeling function that attributes propositions $p\in P$ to the states.
-$\mathcal{L}$ corresponds to the used CPM.
-Figure 2: We need temporary propositions to specify certain properties where the properties are related to labels on the transitions and we do not have a corresponding proposition in the state for them.
-Here is a general example for a temporary proposition, stemming from a transition output.
-Left is the original state ($q$), already labeled with a proposition (AUTHAUTH).
-On the right side, we introduce an internal state ($\tau$) to gain a checkable proposition for a successful read (READOKREADOK).
-It has this temporary proposition and inherits all other propositions from $q$.
+Since we still maintain the Mealy machine labels of transitions, this resulting annotated Mealy machine (AMM) can be seen as combined Mealy-Kripke structure, effectively extending the classical Mealy machine definition with a set of propositions $P$ and a labeling function $\mathcal{L}$.
+Definition 8 (Annotated Mealy Machine).
+We define an annotated Mealy machine (AMM) as $\mathcal{MK}:=(Q,\allowbreak q_{0},\allowbreak\Sigma,\allowbreak\Omega,\allowbreak,\delta,\allowbreak\lambda,\allowbreak P,\allowbreak\mathcal{L})$, with $Q$ being the set of states, $q_{0}$ the initial state, $\Sigma$ the input alphabet, $\Omega$ the output alphabet, $\delta$ the transition function and $\lambda$ the output function of the Mealy machine.
+Finally, $P$ denotes the set of propositions and $\mathcal{L}$ the labeling function that attributes propositions $p\in P$ to the states.
+Figure shows an example AMM that has been generated using the using example Mealy machine (Section 4.1) and CPM.
+We thereby assume that the resource in the Mealy machine is protected, hence all states get the $PROT$ proposition.
+We
+Figure 5: A graphical AMM representation of an example for a simple authenticated file read mechanism.
+The AMM’s proposition set is implicitly defined as the set of all propositions that occur in the gain conditions $C_{g}$ of the proposition map of a corresponding CPM.
+For building the set we can ignore $C_{l}$, because if it contains an additional proposition, it is only subtracted and will therefore not show up in the resulting AMM.
+We can also ignore $C_{\tau}$, as it only contains the implicit state propositions that are also not visible.
+Definition 9 (Annotated Mealy Proposition Set).
+We define the proposition set $P$ of an annotated Mealy machine as the set of all $p$ that occur in one of the gain conditions of a corresponding CPM.
+$P:=\{p\mid\exists(S_{p},S_{\sigma},S_{\omega})\in C_{g};p\in S_{p}\}$
+$\mathcal{L}$ uses the corresponding CPM’s functions.
+We explicitly emphasize that $\mathcal{L}$ resembles the labeling function of a Kripke structure (as it labels states) not an LTS’ labeling function that in contrast labels transitions.
+The latter part is handled by the output function $\lambda$ like in a standard Mealy machine.
+Therefore, CPMs enable to check different complex protocols with the same generic properties, as they instantiate the properties by populating it with the necessary (but protocol-specific) propositions.
+Formally, that means that we execute the gain, loss and $\tau$ functions of a proposition map $C$.
+Definition 10 (Annotated Mealy Labeling Function).
+We define the labeling function $\mathcal{L}$ of an annotated Mealy machine as the sequential execution of the gaingain, lossloss, and $\tau$ functions of an associated Context-based Proposition Map $C$.
+$\mathcal{L}(C,q,\sigma):=\allowbreak{gain}_{c}(q,\sigma)\allowbreak\oplus{loss}_{c}(q,\sigma)\allowbreak\oplus\tau_{c}(q,\sigma)$
+Figure 6: A graphical AMM representation of an example for a simple authenticated file read mechanism with visible implicit states.
+5.1 Annotating Implicit State Propositions
+We only use the implicit state propositions from $C_{\tau}$ in the translation to Rebeca code (Section 6), they are not directly visible in the AMM.
+For any input and output matching $c\in C_{\tau}$, we implement the split translation function (Definition 5) by splitting the respective transition (labeled IN/OUT) into two parts: an implicit transition (labeled IN/OUT) from the origin state ($q$) to a new implicit ($\tau$) state and an empty-input transition (labeled $\varepsilon$$\varepsilon$) from $\tau$ the target state ($q\prime$, which may or may not be equal to $q$).
+The generated $\tau$ state does not have any transitions than these two fragments of the original one.
+We then let state $\tau$ inherit all propositions from $q$ and additionally set the proposition $p$ that is defined in $c$, which is then (only) true in $\tau$.
+Since we do not alter the target state, there is automatically no further inheritance.
+To demonstrate this on our illustrative example, every transition gets an implicit ($\tau$) state, that gets all propositions from the origin state.
+Additionally, the self-transition of $S1$ labeled $READ/OK$ gets the $ACCESSOK$ proposition, since this condition is stated in the example CPM.
+Figure 6 shows a representation of that machine with visible implicit states.
+6 Generating Rebeca Code
+After annotating the Mealy machine with propositions, a fully automated procedure implemented in Java translates the resulting AMM to Rebeca code, where it can be model-checked.
+This section (formally) describes a Rebeca template and the transformation process.
+6.1 Definition of a Rebeca Template
+In the AMM we see input and outputs of each transition.
+In the corresponding Rebeca code we have two actors: environment (an external entity giving input and collecting output) and a system (the entity whose behavior is defined by the AMM).
+Definition 11 (Rebeca Template).
+We define a Rebeca template $\mathcal{R}$ as the parallel execution of a system actor $r_{sys}$ that obtains the annotated Mealy machine’s behavior and an environment actor $r_{env}$ that serves as an input giver.
+$\mathcal{R}:=r_{sys}||r_{env}$
+The system actor $r_{sys}$ holds a state variable with an identifier of the current state and one variable for each proposition from the AMM’s set of propositions (including implicit state propositions).The only known actor is the environment and the method identifiers are equal the set of input symbols plus an additional request method that serves as a calling and reset point.
+Definition 12 (System Actor).
+We define the System Actor $r_{sys}$ as an actor with a state variable as an identifier of the current state ($q\in Q$) from the annotated Mealy machine, initially set to $q_{0}$), each one boolean state variable for the annotated Mealy machine’s set of propositions $P$, and each one boolean state variable for the implicit state propositions in the context-based proposition map.
+Its known actor is the environment actor $r_{env}$.
+It possess a method $f_{\sigma}$ for each input Symbol of the annotated Mealy machine $\sigma\in\Sigma$ and a request method that handles the calls of the system and resets the implicit state propositions.
+$r_{sys}:=\langle V_{sys},M_{sys},K_{sys}\rangle$, with the system variables $V_{sys}:=\{q\}\cup P\cup\{p\mid\exists(S_{p},S_{\sigma},S_{\omega})\in C_{\tau}:p\in S_{p}\}$, the system methods $M_{sys}:=\{f_{\sigma}\mid\sigma\in\Sigma\}\cup\{req()\}$, and the environment as known actor $K_{sys}:=\{r_{env}\}$.
+The environment actor contains the system (Definition 12) as the only other known actor. The set of method identifiers consists the output alphabet.
+Definition 13 (Environment Actor).
+We define the environment actor $r_{env}$ as an actor with the system as known actor ($r_{sys}$).
+The set of method identifiers consists the output alphabet ($f_{\omega}\mid\omega\in\Omega$).
+$r_{env}:=\langle V_{env},M_{env},$ $K_{env}\rangle$), with the implicit state propositions as variable set $V_{env}:=\{\}$ , the environment methods $M_{env}:=\{f_{\omega}\mid\omega\in\Omega\}$ and the system as known actor ($K_{env}:=\{r_{sys}\}$).
+6.2 Template Semantics
+The main part of the Mealy machine’s logic is modeled in the system actor, while the environment actor can be seen as an external input giver to the system.
+To start a Rebeca model built with the defined template, environment’s constructor starts the request message server on the system.
+There, we reset all state variables (which represent the implicit state propositions from $C_{\tau}$) to false and non-deterministically send messages corresponding to each possible input symbol by calling a method $f_{\sigma}$ from $M_{sys}$ via sending the message $\langle r_{sys},r_{sys},f_{\sigma}\rangle$, with a randomly picked $\sigma\in\Sigma$.
+Therefore, a local variable data gets a random value from 0 to $|M_{sys}|=|\Sigma|$ and a switch/case calls one distinct methods from $M_{sys}$.
+In each of the methods of $M_{sys}$ ($f_{\sigma}\in M_{sys}$) four things happen: a) we set the current state identifier ($q\in V_{sys}$) according to the AMM’s transition function $\delta(q,\sigma)$, b) we set all property variables ($p\in V_{sys}$) according to the labeling function ($\mathcal{L}$) in the AMM for the target state [^1] Note: the Method $f_{\sigma}\in M_{sys}$ has a direct, bijective representation in the set of Input Symbols ($\sigma\in\Sigma$): $M_{sys}\leftrightarrow\Sigma$.
+Additionally having the state variable, we can therefore directly use the transition ($\delta$) and output ($\lambda$) functions of the AMM., c) we set a state variable corresponding to a implicit state proposition to true, if there is a respective entry in the CPM ($C_{\tau}$), d) we call a message server of the environment actor from $M_{env}$ ($f_{\omega}\in M_{env}$) that corresponds to the respective output in the respective current state determined by AMM’s output function $\omega=\lambda(q,\sigma)$.
+This means sending the message $\langle r_{sys},r_{env},f_{\omega}\rangle$.
+Each of these output message servers $f_{\omega}\in M_{env}$ (mapping to a $\omega\in\Omega$ from the AMM) calls the request method (with the message $\langle r_{env},r_{sys},{req}\rangle$) to start another interaction cycle.
+This includes, again to reset all implicit state propositions (i.e., resetting the respective state variables in system to $\bot$).
+This assures that each implicit state proposition is present in a state for the model checker, but not maintained further, as it is not part of any state of the AMM.
+6.3 Rebeca Code Transformation
+To transform the AMM into Rebeca code according to the template, we use an automated translation process (which we implemented in Java) that involves basically four steps before writing the actual code.
+Step 1 is parses the DOT file of the machine line by line and create processable data structure for the containing states and transitions, as well as identifying the initial state.
+Step 2 determines for both actors the variables ($V_{sys},V_{env}$) and known counterparts ($K_{sys},K_{env}$).
+Since the template involves two communicating actors, the counterparts are trivial, as they have to simply know each other.
+The state variables of the system actor ($V_{sys}$) are determined by the propositions (one boolean for each $p\in P$), plus an integer to keep track of the system state.
+The environment’s variables ($V_{env})$) consist of all implicit state propositions (every $p$ that occurs in one of the propositions sets in the implicit conditions ($P_{\tau}\in 2^{P}$).
+Step 3 creates the constructor and the request methods of the environment actor.
+The request method (reqreq) is the driving function and randomly calls an input method from the system actor ($f_{\sigma}$) and resets all implicit state proposition variables ($p\in V_{env}$) to false, while the constructor simply calls request to start the model.
+Step 4 parses the transition structure (from Step 1) adds a method for each input and output encountered: an input method $f_{\sigma}$ into the system actor $V_{sys}$ for every input $\sigma\in\Sigma$ and an output method $f_{\omega}$ into the environment actor $V_{env}$ for every output $\omega\in\Omega$.
+Each input method contains a conditional on the current state $q$ and depending on that then a) sets the state variable to $q\prime$ of the transition, b) checks for a proposition delta between $q$ and $q\prime$ and modifies all $p\in V_{sys}$ accordingly and c) calls the output method $f_{\omega}$ that matches $\omega$ in the transition.
+Each output method contains a call to reqreq at the end.
+This implements the implicit transition split described in Section 4: The calls of $f_{\sigma}$ and $f_{\omega}$ represent the first part (transition $q\xrightarrow[\omega]{\sigma}q_{\tau}$, setting the implicit state propositions), while internal call of reqreq from each $f_{\omega}$ represents the second part (transition $q_{\tau}\xrightarrow[\varepsilon]{\varepsilon}q\prime$, resetting the implicit state propositions).
+Further, if a condition for a implicit state proposition matches the output that needs to be distinguished by the input, a parameter for the caller ID (IDσ) is added along with a switch for that parameter, that sets the variable for the implicit state proposition ($(p_{\tau}\in V_{env})\leftarrow\top$).
+This ID is the same the reqreq methods uses to call the respective $f_{\sigma}\in M_{sys}$.
+The ID parameter and switch is omitted if there is less than two options for a calling $f_{\sigma}$ or the lack of matching implicit state propositions.
+Algorithm 2 lists a pseudocode that transforms an AMM ($\mathcal{MK}$, see Definition 8) into Rebeca code according to the given template ($\mathcal{R}$), see Definition 11.
+Algorithm 2 Annotated Mealy Machine to Rebeca Code Transformation
+1://Step 1: Parse AMM
+2:for all lines $\ell$ in DOT file of the AMM do
+3:  if $\ell$ denotes state then
+4:   $Q\leftarrow Q\cup(q\text{ from }\ell),P(q)\leftarrow(P\text{ from }\ell)$
+5:   if $\ell$ declares initial state then
+6:     $q_{0}\leftarrow$ ($q$ from $\ell$)
+7:   end if
+8:  end if
+9:  if $\ell$ denotes transition then
+10:   $t\leftarrow$ ($(q,q\prime,\sigma,\omega)$ from $\ell$); $T\leftarrow T\cup t$
+11:  end if
+12:end for
+13://Step 2: Create Variables and Known Actors
 […]
-5.2 Mapping the Annotated Mealy Machine into Rebeca Code
-After annotating the Mealy machine with propositions, a fully automated procedure implemented in Java translates the resulting annotated Mealy machine to Rebeca code (see a formal definition in Section 5.3).
-In the annotated Mealy machine we see input and outputs of each transition.
-In the corresponding Rebeca code we have two actors: environment (an external entity giving input and collecting output) and a system (the entity whose behavior is defined by the Annotated Mealy machine).
-The environment actor contains a request message server and message servers corresponding to each output in the annotated Mealy machine.
-In the request message server we non-deterministically send messages corresponding to each possible input symbol to the system.
-For these messages related to each input symbol there is message server in the system actor.
-Each of these message servers calls back the message server of the environment actor that corresponds to the respective output in the respective state (by call back we mean sending a return message).
-We keep track of each state using a state variable in the system.
-Additionally, the system entertains state variables for the propositions.
-Every time a proposition changes from one state to another (gaining or losing) it is set to true or false, respectively, along with changing the state and ?
-calling the appropriate output message server on the environment actor.
-Additionally, the algorithm considers the temporary propositions.
-If $C_{\tau}$ defines a temporary proposition, a corresponding state variable is set on the environment actor.
-However, all temporary propositions are reset (i.e., set to false) in the environment actor’s request message server, which is subsequently called by the output message server.
-This assures that the temporary proposition is present in a state for the model checker, but not maintained further, as it is not part of any state of the Mealy machine.
-As a summary, the main part of the Mealy machine’s logic is modeled in the system actor, while the environment actor can be seen as an external input giver to the system.
-5.3 Formal Definition of a Rebeca Template
-We formalize the translation of the annotated Mealy machine by defining a formal Rebeca template by substituting the annotated Mealy components into a Rebeca model definition (see Section 2.4).
-We define this Rebeca model with two actors: $\mathcal{M}:=r_{sys}||r_{env}$$r_{sys}$ is for a system actor that obtains the annotated Mealy machine’s behavior and $r_{env}$ for an environment actor that serves as an input giver.
-We define the system as $r_{sys}:=\langle V_{sys},M_{sys},K_{sys}\rangle$.
-We further define $V_{sys}:=\{q\}\cup AP$.
-So $r_{sys}$ holds a state variable with an identifier of the current state ($q\in Q$ from the annotated Mealy machine, initially set to $q_{0}$) and $P$, which is the annotated Mealy machine’s set of atomic propositions, for each of which $V_{:}{sys}$ contains a boolean state variable.
-The only known actor is the environment ($K_{sys}:=\{r_{env}\}$) and the method identifiers are equal the set of input symbols ($M_{sys}:=\Sigma$).
-The environment ($r_{env}:=\langle V_{env},M_{env},K_{env}\rangle$) contains the temporary propositions from the CPM as state variables ($V_{env}:=C_{\tau})$) and system as other known actor ($K_{env}:=\{r_{sys}\}$).
-The set of method identifiers consists the output alphabet and a request method that handles the calls of the system and resets the temporary propositions ($M_{env}:=\Omega\cup\{req\}$).
-In each of the methods of $r_{sys}$ ($i\in M_{sys}$) two things happen: a) we set the curent state identifier ($q$) along with all properties ($p\in AP$) according to the respective transition in the annotated Mealy machine’s transition function that corresponds to the current state and input symbol $\delta(q,\sigma)$ [^11] Note: the Method $m\in M_{sys}$ has a direct, bijective representation in the set of Input Symbols ($\sigma\in\Sigma$): $M_{sys}\leftrightarrow\Sigma$.
-Additionally having the state variable, we can therefore directly use the transition ($\delta$) and output ($\lambda$) functions of the annotated Mealy machine..
-b) we call an output method in $r_{env}$ ($oinM_{env}$) according to the current sate and the annotated Mealy machine’s output function $\lambda(q,\sigma)$, which means sending the message $\langle r_{sys},r_{env},M_{env}(\omega)\rangle$.
-The environment actor’s ($r_{env}$) request method (reqreq), resets all state variables (which represent the temporary propositions from $C_{\tau}$) to false and randomly picks one input by calling a method from $M_{sys}$ by sending the message $\langle r_{env},r_{sys},M_{sys}(\sigma)\rangle$, with a randomly picked $\sigma\in\Sigma$.
-Each of the other methods (mapping to a $\omega$ from the annotated Mealy machine) sets a state variable corresponding to a temporary proposition to true, if there is a respective entry in the CPM ($C_{\tau}$) [^12] Depening on the calling methods, which is submitted as a parameter in the message.
-and then calls the request method (with the message $\langle r_{env},r_{env},{req}\rangle$) to start another interaction cycle.
-Lastly, we note that the request method is also called by $r_{env}$’constructor, which starts the execution of the Rebeca model ($\mathcal{M}$.
-5.4 Illustrative Example
 […]
-Figure 3: A simple annotated Mealy machine example (left) with an UML diagram of its Rebeca representation (right), as outlined in Section 5.2.
-The black arrows in the UML diagram show an example flow an input of $\sigma_{1}$ in state $q_{1}$, the gray ones a flow of an input of $\sigma_{1}$ in state $q_{2}$.
-Listing 1: Rebeca code from the illustrative example from Section 5.2.
+[…]
+[…]
+18://Step 3: Create environment constructor and system actor’s request method
+[…]
+20:  SYS.REQ()
+21:end method
+[…]
+23:  $\forall p_{\tau}\in V_{sys}:p_{\tau}\leftarrow\bot$
+24:  data $\leftarrow(?\in\{0..(\lvert\Sigma\rvert-1)\})$
+25:  switch data
+26:    case 0: $f_{\sigma_{0}}$() $\in M_{sys}$
+27:    case : …
+28:    case $\lvert\Sigma\rvert-1$: $f_{\sigma_{\lvert\Sigma\rvert-1}}$() $\in M_{sys}$
+29:  end switch
+30:end method
+[…]
+[…]
+33://Step 4: Parse Transitions and Create Methods
+34:for all $t\in T$ do
+35:  method $f_{\omega}$() $\triangleright$ Output functions in $M_{env}$
+36:   req()
+37:  end method
+38:  $M_{env}\leftarrow M_{env}\cup f_{\omega}(\text{})$
+39:  method $f_{\sigma}$() $\triangleright$ Input functions in $M_{sys}$
+40:   for all $s\in Q$ do
+41:     if $s=t[q]$ then
+42:      $V_{sys}[q]\leftarrow t[q\prime]$
+43:      for all $p\in P$ do
+44:        if $q[p]\neq q\prime[p]$ then
+45:         $p\leftarrow q\prime[p]$
+46:        end if
+47:      end for
+48:      if match($c_{\tau}\in C_{\tau}$, t) then
+49:        $V_{sys}[p_{\tau}]\leftarrow\top$
+50:      end if
+51:      $f_{\omega_{t}}$()
+52:     end if
+53:   end for
+54:  end method
+55:  $M_{sys}\leftarrow M_{sys}\cup f_{\sigma}$
+56:end for
+[…]
+Figure 7: An UML sequence diagram of its Rebeca representation (right), as outlined in Section 6.
+Listing 1: Rebeca code from the illustrative example from Section 6.
 reactiveclass ENVIRONMENT(3) {
-statevars {
-boolean omega2set;
 knownrebecs {
 SYSTEM system;
-Environment(){
-self.req();
-msgsrv req() {
-omega2set=false;
-int data = (0);
-switch(data) {
-case 0: system.sigma1(); break;
-msgsrv omega1(){
-self.req();
-msgsrv omega2(){
-omega2set=true;
-self.req();
+ENVIRONMENT() { self.start(); }
+void start() { system.req(); }
+msgsrv err(){ system.req(); }
+msgsrv ok(){ system.req(); }
 reactiveclass SYSTEM(3) {
-statevars {
-int state;
-boolean p;
 knownrebecs {
 ENVIRONMENT environment;
-msgsrv sigma1(){
+statevars {
+boolean prot, auth;
+int state;
+boolean error, accessok;
+SYSTEM() {
+prot=false;
+auth=false;
+state=0;
+msgsrv req() {
+error=false;
+accessok=false;
+int data =?(0,1);
+switch(data) {
+case 0: self.cred(); break;
+case 1: self.read(); break;
+default: self.ERR();
+msgsrv ERR(){
+error = true;
+msgsrv read(){
 if(state==0) {
-state=1;
-p=true;
-environment.omega1();
+state=0;
+environment.err();
 } else
 if(state==1) {
+state=1;
+accessok=true;
+environment.ok();
+msgsrv cred(){
+if(state==0) {
+auth=true;
+state=1;
+environment.ok();
+} else
+if(state==1) {
+auth=false;
 state=0;
-p=false;
-environment.omega2();
+environment.err();
 main {
@@
 Since in Rebeca a model’s behavior is defined similar to a programming language, altering it is trivial.
-Furthermore, Rebeca possess an operator for non-determinism (the ??
-operator).
+Furthermore, Rebeca possess an operator for non-determinism (the ??-operator).
 On the one hand, we can manually add any arbitrary fault that might be interesting to investigate and examine its impact on keeping the properties.
-On the other hand, this allows for re-introducing some non-deterministic behavior that is abstracted away during the learning process (see Section 2.2.1[52].
-By altering the template mentioned in Section 5.3, we just introduce a timeout server on the sender side and alter each input message server on the receiver to trigger a timeout with a certain probability [^14] Realized with a random (??) pick of an integer.
+On the other hand, this allows for re-introducing some non-deterministic behavior that is abstracted away during the learning process [47].
+For instance, we reintroduced timeouts that have been abstracted, by altering the template mentioned in Section 6.1.
+We just introduce a timeout server on the sender side and alter each input message server on the receiver to trigger a timeout with a certain probability, realized with a random (??) pick of an integer.
 If no timeout occurs, the logic continues.
-Otherwise, we reset the state to initial and call the timeout function on the sender, which sets a temporary proposition for timeout, to allow for checking for it.
+Otherwise, we reset the state to initial and call the timeout function on the sender, which sets a implicit state proposition for timeout, to allow for checking for it.
 Since we can set the timeout probability to an arbitrary value, we can play with different scenarios to be checked.
-We also can convert the resulting Rebeca code back to a Mealy machine, so we possess a model that can contain timeouts and errors, just as the originally learned model would have [^15] If the resulting Rebeca code contains non-determinism, the resulting model is not technically a Mealy machine, but a Mealy-styled non-deterministic state model, for it defines multiple transitions for a given input per state to represent the different possibilities..
+We also can convert the resulting Rebeca code back to a Mealy machine, so we possess a model that can contain timeouts and errors, just as the originally learned model would have.
+If the resulting Rebeca code contains non-determinism, the resulting model is not technically a Mealy machine, but a Mealy-styled non-deterministic state model, for it defines multiple transitions for a given input per state to represent the different possibilities.
 Formally, we alter the Rebeca template by extending $M_{env}$ with a timeout method, that sets a state variable ${timeout}\in V_{env}$ (to perform checks for timeouts if needed) and calls the request method.
-Further, we alter each method $m_{sys}\in M_{sys}$ to randomly (with a settable parameter determining the probability) reset the state identifier $q$ to the initial state $q_{1}$ and call the timeout method ($\langle r_{sys},r_{env},{timeout}\rangle$), instead of the output method ($M_{env}(\omega)$ derived from $\lambda(q,\sigma)$.
+Further, we alter each method $m_{sys}\in M_{sys}$ to randomly (with a settable parameter determining the probability) reset the state identifier $q$ to the initial state $q_{1}$ and call the timeout method ($\langle r_{sys},r_{env},{timeout}\rangle$), instead of the output method ($M_{env}(\omega)$ derived from $\lambda(q,\sigma)$).
 6.5 Verifiying the Code
@@
 Based on a previous work in [30], we then use an algorithm that collapses the state space LTS into a Mealy machine and check the bisimilarity of the converted Rebeca Mealy machine with the originally learned model.
+This an extra optional step to provide assurance that no errors have been introduced in the transformation process.
+7 Checking the Model
 The Rebeca model checking tool (RMC) supports Linear Temporal Logic (LTL) to formulate properties.
-For security checking, these LTL form should describe security attributes.
-6 Checking the Model
+For security checking, these LTL formulas should describe security properties.
+Our approach allows us to use generic security properties, which we can concretize for specific state machines using CPMs.
+A property consists of propositions, which the CPM defines (using semantics from the CPM’s protocol context).
+This results in an AMM that includes the propositions, which the security properties we want to check consist of.
+In this chapter we begin with defining four basic, well-established security requirements (authentication, confidentiality, privilege levels, and key validity) and for each define a corresponding security property that we can check.
+Lastly, we briefly describe how to concretize the properties’ propositions using CPMs in two use cases (eMRTDs and UDS).
+Section 9 subsequently describes the practical application and evaluation in these two use cases.
+7.1 Generic Security Requirements
+In this section, we take some well established security properties [42] that are also aligned with our case studies’ standards documents (ICAO [15] and UDS [18]) and formulate them as checkable requirements in given-when-then [^2] Using given, when, and then as keywords in italics.
+format [5]  [^3] We use the terms MUST, MUST NOT, SHOULD, SHOULD NOT and MAY as defined by the Internet Engineering Task Force (IETF) [26], using the same CAPITALIZED style..
+These requirements use attributes, we later (in Section 7.2) use as propositions [^4] Using small capitals style.
+for defining logic properties.
+This definition was performed manually, however, given the generality of the properties and their automated instantiation via CPMs we demonstrate below, the process can be considered a one-time effort.
+If other (either specific or generic) security properties are needed, we recommend to first state a semi-formalized requirement (e.g., in given-when-then format) analyze the components for attributable propositions and constitute the property to meet the requirement’s semantics.
+Subsequently, the properties can be instantiated by populating the propositions with a concrete system context with a CPM.
+7.1.1 Authentication
+To establish the authenticity as a property, we must require authentication of a user before accessing protected resources (not every resource must be protected, some may be openly accessible – this depends on the system’s security design).
+As a requirement, we define it as follows:
+Requirement R1:Given the system is not in an authenticated state, when an access operation on a protected resource occurs, then the operation MUST NOT return a positive response.
+7.1.2 Confidentiality
+To acquire confidentiality (also called secrecy in cryptographic contexts), most protocols allow for encryption (and may require it under certain conditions).
+We therefore recommend encryption for protected resource access:
+Requirement R2:Given the system is in an arbitrary state, when an unsecured access operation on a protected resource occurs, then it SHOULD NOT return a positive response.
+unsecured access thereby denotes a file operation without cryptographic security controls (i.e., encryption).
+7.1.3 Privilege Levels
+It is common practice for more complex systems and protocols to define more than one privilege level according to different roles and corresponding rights of an actor in order to provide access to more critical resources.
+This also applies for UDS, where this is implemented by different session and security levels [19].
+Likewise, for eMRTDs, sensitve data have to be additionally protected via terminal authentication.
+Requirement R3:Given the level of privilege is not sufficient, when an access operation on a critical resource occurs, then the operation MUST NOT be successful.
+A sufficient level of privileges also implies authentication.
+7.1.4 Key Validity
+Another important property of secure connections is key validity and recency (i.e., no newer key must exist).
+Since we do not consider a time component in the automata, we do not use recency in the sense of freshness.
+An invalid key in this sense is either a wrong key or an old key (not recent).
+Requirement R4:Given any operation that requires a key, when an invalid key is provided, then the key MUST NOT be accepted.
 7.2 Defining Generic Properties
-Since the requirements in Section 3 are very generic, we can also define generic properties to check them.
-The become more specific in conjunction with the CPMs (Section 5.1), which define the meaning of propositions used in the property in the context of a specific protocol [^16] If a proposition used in a property is not defined in the respective CPM, we define it as false($\bot$) for the respective property..
-For the sake of readability, we restrict our definitions of access to read operations but write, update, and similar operations can be defined the same way.
+Since the requirements in Section 7.1 are generic, we can also define generic properties to check them.
+Generally, we use the given and when parts as a premise and the then part as a consequence (except for the mapping from Requirement R7.1.4 to property P7.2.4, for which we trivially use no invalid keys as a property).
+These become more specific in conjunction with the CPMs, which define the meaning of propositions used in the property in the context of a specific protocol.
+If a proposition used in a property is not defined in the respective CPM, we have two options: a) leave it undefined, which creates an error in the model checking and makes the user aware of this fact or b) define it as false($\bot$) for the respective property.
+The latter option means that, for semantically sensible settings, our properties hold with undefined propositions, effectively ignoring the concerned proposition.
+For the sake of readability, we restrict our definitions of access to read operations; write, update, and similar operations can be defined the same way.
 7.2.1 Authentication
-For Requirement R3.1, we define an authenticated state as a state with the proposition authenticated set to false, a when an access operation on a protected resource occurs, then the operation must not return a positive response (access ok).
+For Requirement R7.1.1, we define an authenticated state as a state with the proposition authenticated set to true.
+If this is not the case, when an access operation on a protected resource occurs, then the operation must not return a positive response.
 Property P1: $\square(\neg{AUTH}\land{PROT}\rightarrow\neg{ACCESSOK})$ [^5] Afra does not support implications in the property file, therefore we use $\square(\neg(\neg{AUTH}\land{PROT})\lor\neg{ACCESSOK})$, instead..
-Here, we define AUTH{AUTH} for authenticated, PROT{PROT} for protected, and ACCESSOK{ACCESSOK} for a positive response [^18] Example for non-existing protocols: if we have a protocol with no protected resources, we set $PROT:=\bot$, which means that reading is always allowed according to Property P6.1.1.
-(i.e., a successful read operation).
+Here, we define AUTH{AUTH} for authenticated, PROT{PROT} for protected, and ACCESSOK{ACCESSOK} for a positive response (i.e., a successful read operation).
+If we have a protocol with no protected resources, we can set $PROT:=\bot$, which means that reading is always allowed according to Property P7.2.1.
 7.2.2 Confidentiality
-For checking the confidentiality (Requirement R3.2), we assume the presence of unsecured and secured operations.
+For checking the confidentiality (Requirement R7.1.2), we assume the presence of unsecured and secured operations.
 For the access operation as used in Section 7.2.1, we therefore define two exemplary subsets: an unsecured read $UREAD$ and a secured read $SREAD$.
-Since $READ$ is a superset, in the respective context maps we then define that $READ$ will always be set along with $UREAD$ or $SREAD$ is set (using the same conditions, i.e., $c_{READ}\leftrightarrow c_{UREAD}\lor c_{SREAD}$).
-Given requirement R3.2, we define that an $UREAD$ may not succeed when $PROT$ is true.
-Please note that we fail to apply this property to the UDS case, since there is no encrypted communication on the CAN bus that servers as a medium for the UDS protocol [^19] While there are some proprietary and research solutions for encrypted CAN, none of them is widely adopted (i.e., used by a larger number of major OEMs).
-An industry standard gaining momentum, the AUTOSAR Secure On-board Communication (SecOC) [54], only provides authentication and integrity, but not confidentiality..
+Since $READ$ is a superset, in the respective context maps we then define that $READ$ will always be set along with $UREAD$ or $SREAD$ is set (using the same conditions, i.e., $c_{READ}=c_{UREAD}\lor c_{SREAD}$).
+Given requirement R7.1.2, we define that an $UREAD$ may not succeed when $PROT$ is true.
+Please note that we fail to apply this property to the UDS case, since there is no encrypted communication on the CAN bus that servers as a medium for the UDS protocol.
+While there are some proprietary and research solutions for encrypted CAN, none of them is widely adopted (i.e., used by a larger number of major OEMs).
+An industry standard gaining momentum, the AUTOSAR Secure On-board Communication (SecOC) [3], only provides authentication and integrity, but not confidentiality.
 Property P2: $\square({PROT}\rightarrow\neg\text{ }{UREADOK})$
 7.2.3 Privilege Levels
-We assume that there are areas that go beyond a basic protection level and that there additional protection measures (e.g., different type or level of authentication) have to be taken (Requirement R3.3).
+We assume that there are areas that go beyond a basic protection level and that there additional protection measures (e.g., different type or level of authentication) have to be taken (Requirement R7.1.3).
 We define this higher protection level as critical resources ($CRIT$) and the additional authentication as privileged ($PRIV$).
@@
 7.2.4 Key Validity
-A property for using a valid key (Requirement R3.4) is straightforward, since it only has to define that an invalid key must not be accepted in the entire system.
+A property for using a valid key (Requirement R7.1.4) is straightforward, since it only has to define that an invalid key must not be accepted in the entire system.
 Property P4: $\square(\neg{INVKEYOK})$
@@
 7.3 Checking the Properties
-This section defines and explains the CPMs for eMRTDs and UDS, along with defining some protocol-specific properties that allow for checking the protocol flow.
-6.2.1 eMRTDs
-For eMRTDs, we define $PROT$ as $DF$ input in $C_{g}$ of the CPM for eMRTDs (Table 1), since the ICAO standard defines the applications to be in a protected zone [^20] Note: we only set DF for selecting LDS1, which corresponds to the eMRTD application.
+Once we have a Rebeca model according to a template (Definition 11) and respective properties, the Rebeca Model Checker (RMC) can check the model for these properties.
+Using a protocol-specific CPM, we define the propositions that constitute the properties in the context of a specific protocol.
+In our running example (see Figure 6), we defined the necessary propositions for Property P7.2.1 only (for the sake of simplicity), which holds for this example.
+If we alter the transition for $\sigma=READ$ in $q=s0$ from $\omega=ERR$ to $\omega=OK$, then respective implicit ($\tau$) state gets the proposition $ACCESSOK$, but not $AUTH$ and the property will be violated.
+Semantically the latter case would mean a successful access without authentication.
+7.3.1 Checking eMRTDs
+For eMRTDs, we define $PROT$ as $DF$ input in $C_{g}$ of the CPM for eMRTDs (Table 3), since the ICAO standard defines the applications to be in a protected zone.
+We only set DF for selecting LDS1, which corresponds to the eMRTD application.
 Technically, DF would also be set when selecting other applications (i.e., LDS2).
 However, we then would design the CPM to distinguish between DFs.
-This is not necessary in our case, since we only use one., while $AUTH$ is gained by performing a basic authentication (BAC) operation.
+This is not necessary in our case, since we only use one.
+$AUTH$ is gained by performing a basic authentication (BAC) operation.
 As $ACCESSOK$ we define a successfully performed file selection in $C_{\tau}$.
-This means that for P6.1.1 a BAC must be performed before selecting any file inside an application.
-For $UREADOK$, we define an unsecured read binary operation. This means that for P6.1.2, only secured read binary operations are allowed.
+This means that for P7.2.1 a BAC must be performed before selecting any file inside an application.
+For $UREADOK$, we define an unsecured read binary operation. This means that for P7.2.2, only secured read binary operations are allowed.
 We define $CRIT$ as a successful selection of data groups 2 or 3 in $C_{\tau}$.
@@
 Since, we do not have an additional authentication method implemented, $PRIV$ is set to false ($\bot$).
-This means that P6.1.3 only holds when these sensitive data groups cannot be selected in the complete model.
-Lastly, we define $INVKEYOK$ for P6.1.4 as any successful secured operation that has been carried out using an old key or an all-zero key in $C_{\tau}$.
-We realize these by dedicated input symbols (see Section 4.1.1).
+This means that P7.2.3 only holds when these sensitive data groups cannot be selected in the complete model.
+Lastly, we define $INVKEYOK$ for P7.2.4 as any successful secured operation that has been carried out using an old key or an all-zero key in $C_{\tau}$.
+We realize these by dedicated input symbols starting with WS (wrong key secured) and OS (old key secured).
 We also use a couple of complimentary properties for checking eMRTDs.
@@
 This also creates the need for a normal read operation not to work on protected resources: PlainRead: $\square(\neg(UREADOK\land(\neg EF\lor DF)))$, which reads as a successful plain read (READOK) cannot happen without a selected resource outside of LDS1 ($\neg EF$$\lor DF$).
-We can further specify a secure select process with the following property: SecureReadFollowsSecureSelect: $((\neg SREADOK)\ \mathcal{U}\$$\text{SSELEFOK})\lor\square(\neg SREADOK)$.
+We can further specify a secure select process with the following property: SecureReadFollowsSecureSelect: $((\neg SREADOK)\ \mathcal{U}\ \allowbreak\text{SSELEFOK})\allowbreak\lor\square(\neg SREADOK)$.
+This means that a secured read operation may only follow after a secured selection operation has occurred, independent of the EF/DF properties in the AMM.
+The reason is to prevent errors like that the EF proposition has been obtained through an unsecured select EF.
 Table 3 gives an overview of the CPM.
-6.2.2 UDS
+7.3.2 Checking UDS
 Authentication in UDS works over SecurityAccess and subsequent SecurityAccessWithKey.
-We therfore define the $AUTH$ proposition after the second step has been successfully exectued.
-As a canary for P6.1.1, we perform a reading operation on a protected resource.
+We therefore define the $AUTH$ proposition after the second step has been successfully executed.
+To verify property P7.2.1, we perform a reading operation on a protected resource.
 We choose to start the Check Programming Dependencies (0xFF01) routine.
-This routine does not alter the firmware and is standardized [8].
-Furthermore, it is industry practice that this route is protected by a Security Access routine [55].
+This routine is standardized [19] and does not alter the firmware.
+Yet, it is industry practice that this route is protected by a Security Access routine as it is a preparatory action for flashing the system [56].
 Therefore we set both $PROT$ and $ACCESSOK$ as a $c_{\tau}$ when starting this routine.
-For P6.1.2, set perform a direct memory read operation to set $UREADOK$.
+For P7.2.2, set perform a direct memory read operation to set $UREADOK$.
 Unfortunately, we don not have the knowledge of which memory are actually protected (as a workaround defining the CheckASWBit routine could also be defined as $PROT$).
-For P6.1.3, we define the RequestDownload routine as $CRIT$, as it allows to actually flash the device, while $PRIV$ his defined as SecurityAcccess with a higher level.
-Lastly for P6.1.4, $INVKEYOK$ is defined as SecurityAcccess with a wrong key (as an own input symbol).
+For P7.2.3, we define the RequestDownload routine as $CRIT$, as it allows to actually flash the device, while $PRIV$ his defined as SecurityAcccess with a higher level.
+Lastly for P7.2.4, $INVKEYOK$ is defined as SecurityAcccess with a wrong key (as an own input symbol).
 Table 4 shows the CPM for UDS.
@@
 Otherwise, the trace can be fed back to the learning system to refine the model.
+We implement this step using a custom Oracle for the automata learner, where we can input the found counterexample for a new learning iteration.
+As a by-product, we use this custom oracle to improve the learning performance with a priori protocol knowledge by using test traces that resemble the standard behavior of a protocol first.
 The verification of found traces is trivial, as the same tool set from the learning can be used for concretizing the input and interfacing with the SUL.
@@
 CRIT EF* 9000
-Temporary ($C_{\tau}$)
+Implicit State Propositions ($C_{\tau}$)
 UACCESSOK, ACCESSOK SEL_EF* 9000
@@
 Though the use cases are very different, in both cases the model checking is very similar.
-At the present abstraction level, we check generic security requirements (Section 3) using according LTL properties (Section 6.1).
+At the present abstraction level, we check generic security requirements using according LTL properties.
+The properties and CPMs are pre-defined (whereby the specification is fairly easy), while the annotated Mealy is automatically generated (utilizing a CPM) from a fully automatically learned Mealy machine.
+Thus, we can claim that the complete process of learning and checking a real-life eMRTD and UDS-based system is largely automated, with very little human interaction necessary.
 9.1 Electronically Machine-Readable Travel Document
-For the eMRTD use case, we scrutinize two different Austrian passports, on elder, expired and one current (see a process based on equivalence checking in [34]).
+For the eMRTD use case, we scrutinize two different Austrian passports, one elder, expired and one current (see a process based on equivalence checking in [29]).
 The model was learned via LearnLib using the TTT algorithm (with binary closure) and a minimum input trace length of 40 symbols, a maximum of 50, and 150 random walk conformance tests as equivalence oracle.
-Figure 4: Simplified version of the annotated eMRDT.
+Figure 8: Simplified version of the learned eMRDT’s AMM.
 Self-loops that do not add to the understanding have been removed for readability.
+PP_ is a general prefix for passport inputs.
 (S)SEL_EF* (ex) and (noex) denotes all elementary file selections of existing and non-existing files.
-Using the method described above, we translated the Mealy Machine into Rebeca code (see Listings LABEL:lst:NFCsend and LABEL:lst:NFCrec for parts of the model code for the environment and system agent, respectively).
-Using the CPM and properties as described in Section 6.2.1, we were able to verify that described security properties hold the scrutinized systems.
+O—WS* denote secure commands that are incrypted with a wrong or old key.
+Note that with our CPM definition for eMRTDs, the protocted attribute (PROT) is equal to a selected application (DF).
+The result was a 6 state Mealy machine which we annotated the Mealy machine to an AMM using the method described above.
+The AMM (see Figure 8) shows how files are selected (EF) inside and outside of the eMRTD application (DF/PROT) and the authentication inside the application (AUTH).
+It is noteworthy that a successful file read in this diagram only occurs via SRD_BIN in state S4, which is a visual evidence that P7.2.1 and P7.2.2 hold.
+The higher privileged data groups where not selectable, since we do not have higher privileges (P7.2.3 holds).
+Old and wrong key commands result in de-authentication instead of a success, as can be seen with the transitions of PP_O—WS (denoting old/wrong key operations) with an error in states S3 (falling back to s2) and S4 (falling back to S5), showing that P7.2.4 holds.
+For the automated formal analysis of these properties through model checking, we automatically translated the AMM into Rebeca code (see the Listings in A for parts of the model code for the environment, the system agent, and the property file respectively).
+Using the CPM in Table 3 and properties as described in Section 7.3.1, we were able to verify that described security properties hold the scrutinized systems.
+The testing is step therefore not applicable, since no property violation was present.
+This means that we could verify in an automated pipeline, that Austrian passports implement authentication, confidentiality, privilege levels, and key validity.
+Furthermore, we also successfully verified the additionally defined properties for passports, e.g.
+that a successful secure read may only occur after a successful secure select.
+9.2 Automotive Electronic Control Unit
+Our example use case for UDS is an automotive electronic control unit (ECU) from a major European Tier-1 supplier running in a vehicle from a Chinese car manufacturer.
+The model was learned via LearnLib using the TTT algorithm with a minimum input trace length of 20 symbols, a maximum of 50, and 50 random walks as equivalence oracle.
+We received an 8 state Mealy machine, which was annotated to an AMM (see Figure 9) using the CPM in Table 4) of this ECU.
+Table 4: CPM for UDS (SA means SecurityAccess).
+Proposition Input Output
+Gains ($C_{g}$)
+AUTH SAWithKey 67
+EXT Extended 5003
+PROG Programming 5002
+PRIVa HLSAWithKey 67
+Losses ($C_{l}$)
+EXT,PROG Default 5001
+EXT Programming 5002
+PROG Extended 5003
+AUTH SA, SAwKey, SAwWrongKey 7f
+AUTH Session 50
+Implicit State Propositions ($C_{\tau}$)
+INVKEYOK, WRONGKEYOK SAwWrongKey 67
+PROT CheckASWBit 71
+ACCESSOK CheckASWBit 71
+UACCESSOK CheckASWBit 71
+CRIT RequestDownload 74
+UREADOK Read* 62
+  • a
+    Since, we do not have higher level security access, this condition is hypothetical (will not be triggered).
+Figure 9: Simplified AMM of the learned ECU model.
+Self-loops that do not add to the understanding have been removed and the output truncated for readability.
+The star (*) denotes any other input not explicitly stated.
+The transition in red shows the violation of security requirement R7.1.4 and property P7.2.4.
+Automatically translating the AMM into Rebeca (see the Listings in B for parts of the model code for the environment, the system agent, and the property file respectively), we subsequently use the properties from Section 7.2 to check the model for security properties, along with the CPM mentioned above.
+We could verify P7.2.1.
+For P7.2.2, we lack the knowledge which memory areas are actually protected.
+However, it is well known that the CAN bus (over which UDS communication runs) is completely unencrypted, while principally critical (safety) data can be exchanged [58].
+Therefore, R7.1.2 is not met by the protocol itself, regardless of any implementation.
+To test P7.2.3, we lack a higher layer security access.
+However, scrutinizing critical functions (RequestDownload), we saw P7.2.3 to hold, as it were not possible to trigger with our current security level.
+P7.2.4 was found to be violated.
+In the authenticated state S5 (see the transition in red in Figure 9), the system accepts a wrong key [^6] An issue known to us from previous works [10].
+with a success return code.
+According to the standard, wrong key should not be acknowledged, even though a wrong key does not mean the security level must be locked (de-authentication) [18].
+For the testing step, the learned input trace of S5 and a wrong key input can be applied to the system using the automata learning adapter class.
+We omitted this step, since this is a known issue on this device, so we can be sure this is not a false positive.
+As a summary, we can state that the examined ECU only fulfills two out of four tested security requirements.
+10 Related Work
+This work is based on previous work on model learning and model checking [10, 29, 30].
+Peled et al.
+[37] have provided a very influential paper regarding black box testing that combines automata learning with model checking.
+They learn a model (using Angluin’s algorithm) create a cross product with a Büchi automaton created from the negate of an LTL property to check.
+If the accepted language is empty, the property is satisfied.
+Otherwise, the contained (non-empty) words it serve as a counterexamples that will be tested against a true system model.
+If these tests yield the same results, the counterexample poses a property violation.
+Otherwise, it serves as a counterexample for refining the model.
+They propose both an off-line and on-the-fly approach.
+The former uses a fully learned (i.e., a full run-through of Angluin’s algorithm) model that is checked.
+The latter incorporates the checking into the learning process and stops it, if a violation is found, potentially greatly reducing the runtime.
+Groce et al.
+[14] advance the former approach by integrating the model checking deeper in the learning algorithm for an efficiency boost.
+Shijubo et al [45] expand Peled’s approach by introducing strengthened specifications, that tend to find counterexamples that earlier detect divergence between the learned and the actual model, boosting the learning process’ efficiency.
+We refrain from using on-the-fly checking because we want to fully investigate the model for any violation of a defined set of properties, and therefore not stop at the first found violation, as the purpose is to investigate already implemented real-world systems for all incorrect behavior.
+While the Rebeca model checker also stops with a counterexample when a property violation is found, we can, in contrast to on-the-fly-checking, check the other properties as well on the complete model and therefore determine if the system violates multiple properties.
+Furthermore, many properties do not actually require a modal operator – e.g., Property 7.2.1 uses a single $\square$ operator, which means it should hold globally and a further distinction is not necessary.
+We can check these with a simple script running over a Rebeca-generated full state space, enabling us to find all occurrences of a violation.
+While our work is based on a similar idea as the off-line variant, we extended the approach by annotating Mealy machines with atomic propositions creating AMMs that we turned into Rebeca code for model checking.
+That way, we implemented all the phases for going through learning, checking, and testing real-world examples.
+In our case, however, we automated the complete process by combining (a) the learner(s) (using LearnLib [21]), (b) an automated way of annotating Mealy machines with propositions, creating AMMs and translating it into a checkable model, and (c) a model checker translates this model and (the negation) LTL properties to Büchi automata, creates a product, and check its language for emptiness [23] with generic properties.
+This improves both the usability of the approach and its applicability to real-world systems.
+Neider and Jansen used a symbolic approach to learn and model-check DFAs (but not Mealy or More machines) [34].
+Similarily, Fiterau-Brostean et al.
+[11, 12] also used a symbolic approach (using the NuSMV model checker).
+Translating the learned model to Rebeca, our approach in contrast provides more advanced possibilities to manipulate the model, including re-introducing abstracted (e.g.
+non-deterministic) behavior, as they also stated in their work that ironing out timing issues were a major engineering problem.
+Altering the Rebeca template, it is trivial to re-introduce timeouts (non-deterministically) to check a protocol’s behavior under such conditions.
+This allows for more advanced checking possibilities.
+Furthermore, our approach using CPMs does not only allow for a more explicit and comprehensible annotation of states with propositions, but also using generic properties and a clean separation between state propositions (visible in the annonated Mealy machine) and pure transition outputs (in $C_{\tau}$ and $\tau$ states).
+For the principle conversion of LTS (which Mealy machines could be seen as) and Kripke structures there is a well-known standard procedure by De Nicola and Vaandrager [9], used by many scholarly works .
+However, our approach differs by utilizing both state and transition labels’ information to answer very specific questions in the security context (e.g., about resource access and prior authentication) using implicit state propositions based on the state labels.
+Also, the proposed AMM structure is more convenient to transfer into a Rebeca model, which is beneficial, as stated above.
+Furthermore, we argue that the CPM approach provides an easy access for model checking, as these are easy to create and the rest of the process is automated.
+Other approaches [44, 50, 27] combine automata learning with bisimulation-based equivalence checking, which need a specification model to compare the learned automaton with.
+Our approach needs generalized properties.
+We do not know of an approach that translates a learned model into a modeling language allowing for both direct property checking and manipulating the model to simulate scenarios.
+11 Conclusion
+In this paper, we presented an approach to combine automata learning with model checking for reactive systems.
+We defined Context-based Proposition Maps (CPMs) that provide an annotation mechanism to annotate Mealy machines inferred by Automata Learning with propositions.
+As a result, we received annotated Mealy machines (AMMs) that combine attributes of Mealy machines and Kripke structures.
+We also formally defined these AMMs as an extension of classical Mealy machines combined with a set of properties and a (state) labeling function.
+We subsequently translated the AMMs into Rebeca code in an automated process by formally defining a template using two actors: a receiver modeling the learned system and a sender modeling an external actor that interacts with the system.
+Starting from four high-level security requirements (authentication, confidentiality, privilege levels, and key validity), we defined generic LTL properties to check models for security.
+Protocol-specific CPMs provided the context to assure the generated Rebeca code provides all propositions that are used in these properties.
+This way, each property can be checked with regard to the protocol the examined system runs on.
+We presented a case study with systems from different domains (two passports and an automotive control unit) speaking different protocols (NFC and UDS) to show the versatility of the approach on real-world off-the-shelf systems.
+We were able to verify the security properties for all requirements on the two different passports.
+On the automotive control unit, our verification process verified only two of them (authentication and privilege levels).
+The third (confidentiality) is intrinsically not provided by the protocol.
+The last one (key validity) was not met (property violated) on the examined unit, which means we could falsify this requirement.
+These techniques applied in the presented case studies provide evidence for the usefulness of the approach, since it significantly reduces modeling and property design efforts through automation.
+Given a working adapter for learning a protocol, the only manual effort to be taken by the verifier is to define a corresponding CPM.
+As we outlined in this paper, this task is not particularly resource-intense.
+As a summary, we provided security properties for an automated model checker (RMC) by deriving them from generic security requirements and their assigning their propositions with CPMs.
+On the other hand we automatically derived checkable models by generting Mealy machines through automata learning and annotating them to AMMs using the same CPMs as for the properties.
+11.1 Outlook
+Further research directions to lift the approach to a larger scale lie in creating adapters for other different protocols and creating other generalized LTL properties to check different aspects of security.
+Additionally, we also investigate utilizing an LLM to generate LTL properties from threat models to check their respective implementations’ Mealy machines.
+Also we can improve the learning process by using timing [53] and by extending the model with unknown inputs using input symbol mutation; if the unknown input triggers new behavior, we can dynamically extend the input alphabet (as described here [43]).
+This allows for a more holistic testing.
+Furthermore we plan to extend the CPMs to be more expressive (e.g., allow for logic operators in the CPM conditions or add an operator to assign propositions to the initial state).
+12 Acknowledgement
+This work has received funding from the European Union under the Horizon Europe programme under grant agreement No.
+101168438 (project INTACT) as well as from the Swiss State Secretariat for Education, Research and Innovation and UK Research and Innovation under grant No.
+10126241.
+We further acknowledge the support of the Swedish Knowledge Foundation via the industrial doctoral school RELIANT under grant No.
+20220130.
+The authors want to thank the anonymous reviewers of the Computers & Security Journal, as well as of the International Conference on Runtime Verification for their valuable input.
+We further want to thank Thomas Grandits for helping to formulate requirements and properties more understandable.
+Appendix A eMRTD Rebeca Code Listings
 Listing 2: Example of generated eMRTD environment actor code.
+reactiveclass ENVIRONMENT(3) {
+knownrebecs {
+SYSTEM system;
+ENVIRONMENT() {
+self.start();
+void start() {
+system.req();
+msgsrv req_6986(){
+system.req();
+msgsrv req_9000(){
+system.req();
+Listing 3: Example of generated eMRTD system actor code.
+reactiveclass SYSTEM(3) {
+knownrebecs {
+ENVIRONMENT environment;
+statevars {
+boolean ef, df, prot, auth;
+int state;
+boolean error;
+boolean ureadok, sreadok, readok, sselefok, selefok, accessok, uaccessok, saccessok, invkeyok, wrongkeyok, oldkeyok;
 msgsrv req() {
@@
 case 32: system.pp_ssel_ef_dg1(); break;[..]
-msgsrv req_9000(int data){
-switch(data) {
-case 0: accessok=true;selefok=true;
-uaccessok=true;break;
-case 1: break; [..]
-case 22: readok=true;ureadok=true;break; [..]
-case 28: break; [..]
-case 32: accessok=true;sselefok=true;
-saccessok=true;break; [..]
-self.req();
-Listing 3: Example of generated eMRTD system actor code.
 msgsrv pp_rd_bin(){
@@
 state=5; environment.req_6988();
-8.2 Automotive Electronic Control Unit
-Our example use case for UDS is an automotive electronic control unit (ECU) from a major European Tier-1 supplier running in a vehicle from a Chinese car manufacturer.
-The model was learned via LearnLib using the TTT algorithm with a minimum input trace length of 20 symbols, a maximum of 50, and 50 random walks as equivalence oracle.
-Figure 5 shows the annotated Mealy machine (using the CPM in Table 2) of this ECU.
-Table 2: CPM for UDS (SA means SecurityAccess).
-Proposition Input Output
-Gains ($C_{g}$)
-AUTH SAWithKey 67
-EXT Extended 5003
-PROG Programming 5002
-PRIVa HLSAWithKey 67
-Losses ($C_{l}$)
-EXT,PROG Default 5001
-EXT Programming 5002
-PROG Extended 5003
-AUTH SA, SAwKey, SAwWrongKey 7f
-AUTH Session 50
-Temporary ($C_{\tau}$)
-INVKEYOK, WRONGKEYOK SAwWrongKey 67
-PROT CheckASWBit 71
-ACCESSOK CheckASWBit 71
-UACCESSOK CheckASWBit 71
-CRIT RequestDownload 74
-UREADOK Read* 62
-  • a
-    Since, we do not have higher level security access, this condition is hypothetical (will not be triggered).
-Figure 5: Simplified annotated Mealy machine of the learned ECU model.
-Self-loops that do not add to the understanding have been removed and the output truncated for readability.
-The star (*) denotes any other input not explicitly stated.
-We subsequently use the properties from Section 6.1 to check the model for security properties, along with the CPM for UDS (Section 6.2.2).
-We could verify P6.1.1 .
-For P6.1.2, we lack the knowledge which memory areas are actually protected.
-However, it is well known that the CAN bus (over which UDS communication runs) is completely unencrypted, while principally critical (safety) data can be exchanged [56].
-Therefore, R3.2 is not met by the protocol itself, regardless of any implementation.
-To test P6.1.3, we lack a higher layer security access.
-However, scrutinizing critical functions (RequestDownload), we saw P6.1.3 to hold, as it were not possible to trigger with our current security level.
-P6.1.4 was found to be violated [^21] According to the standard, wrong key should not be acknowledged, even though a wrong key does not mean the security level must be locked (de-authentication) [39]., as in the (authenticated) state S6 (see Figure 5), the system accepts a wrong key [^22] An issue known to us from previous works [48]..
-9 Related Work
-This work is based on previous work on model learning and model checking [48, 34, 53].
-Peled et al.
-[19] have provided a very influential paper regarding black box testing that combines automata learning with model checking.
-They learn a model (using Angluin’s algorithm) create a cross product with a Büchi automaton created from the negate of an LTL property to check.
-If the accepted language is empty, the property is satisfied.
-Otherwise, the contained (non-empty) words it serve as a counterexamples that will be tested against a true system model.
-If these tests yield the same results, the counterexample poses a property violation.
-Otherwise, it serves as a counterexample for refining the model.
-They propose both an off-line and on-the-fly approach.
-The former uses a fully learned (i.e., a full run-through of Angluin’s algorithm) model that is checked.
-The latter incorporates the checking into the learning process and stops it, if a violation is found, potentially greatly reducing the runtime.
-Groce et al.
-[57] advance the former approach by integrating the model checking deeper in the learning algorithm for an efficiency boost.
-Shijubo et al [58] expand Peled’s approach by introducing strengthened specifications, that tend to find counterexamples that earlier detect divergence between the learned and the actual model, boosting the learning process’ efficiency.
-We refrain from using on-the-fly checking because we want to fully investigate the model for any violation of a defined set of properties, and therefore not stop at the first found violation, as the purpose is to investigate already implemented real-world systems for all incorrect behavior.
-[^23] While the Rebeca model checker also stops with a counterexample when a property violation is found, we can, in contrast to on-the-fly-checking, check the other properties as well on the complete model and therefore determine if the system violates multiple properties.
-Furthermore, many properties do not actually require a modal operator – e.g., Property 6.1.1 uses a single $\square$ operator, which means it should hold globally and a further distinction is not necessary.
-We can check these with a simple script running over a Rebeca-generated full state space, enabling us to find all occurrences of a violation.
-While our work is based on a similar idea as the off-line variant, we extended the approach by annotating Mealy machines with atomic propositions creating Kripke-like structures that we turned into Rebeca code for model checking.
-That way, we implemented all the phases for going through learning, checking, and testing real-world examples.
-In our case, however, we automated the complete process by combining (a) the learner(s) (using LearnLib [10]), (b) an automated way of generating Kripke structures and translating it into a checkable model (described in Section 5) and (c) a model checker translates this model and (the negation) LTL properties to Büchi automata, creates a product, and check its language for emptiness [9] with generic properties (Section 6).
-This improves both the usability of the approach and its applicability to real-world (cyber-pyhsical) systems.
-Neider and Jansen used a symbolic approach to learn and model-check DFAs (but not Mealy or More machines) [59].
-Similarily, Fiterau-Brostean et al.
-[60, 61] also used a symbolic approach (using the NuSMV model checker).
-Translating the learned model to Rebeca, our approach in contrast provides more advanced possibilities to manipulate the model, including re-introducing abstracted (e.g.
-non-deterministic) behavior (see Section 5.5), as they also stated in their work that ironing out timing issues were a major engineering problem.
-Altering the Rebeca template, it is trivial to re-introduce timeouts (non-deterministically) to check a protocol’s behavior under such conditions.
-This allows for more advanced checking possibilities.
-Furthermore, our approach using CPMs does not only allow for a more explicit and comprehensible annotation of states with propositions, but also using generic properties and a clean separation between state proposistions (visible in the annonated Mealy machine) and pure transition outputs (in $C_{\tau}$ and $\tau$ states).
-Other approaches [62, 15, 63] combine automata learning with bisimulation-based equivalence checking, which need a specification model to compare the learned automaton with.
-Our approach needs generalized properties.
-We do not know of an approach that translates a learned model into a modeling language allowing for both direct property checking and manipulating the model to simulate scenarios.
-10 Conclusion
-In this paper, we presented an approach to combine automata learning with model checking for cyber-physical systems.
-We defined Context-based Proposition Maps (CPMs) that provide an annotation mechanism to annotate Mealy machines inferred by Automata Learning with propositions.
-As a result, we received annotated Mealy machines that combine attributes of Mealy machines and Kripke structures.
-We subsequently translated the annotated Mealy machine into Rebeca code by formally defining a template using two actors: a receiver modeling the learned system and a sender modeling an external actor that interacts with the system.
-Starting from four high-level security requirements (authentication, confidentiality, privilege levels, and key validity), we defined generic LTL properties to check models for security.
-Protocol-specific CPMs provided the context to assure the generated Rebeca code provides all propositions that are used in these properties.
-This way, each property can be checked with regard to the protocol the examined cyber-physical system runs on.
-We presented a case study with systems from different domains (a passport and and automotive control unit) speaking different protocols (NFC and UDS) to show the versatility of the approach.
-We were able to verify the security properties for all requirements on to different passports.
-On the automotive control unit, we could verify two of them (authentication and privilege levels), while one of them (confidentiality) is intrinsically not provided by the protocol and the last one (key validity) was not met (property violated) on the examined unit.
-10.1 Outlook
-Further research directions to lift the approach to a larger scale lie in creating adapters for other different protocols and creating other generalized LTL properties to check different aspects of security.
-Additionally, we also investigate utilizing an LLM to generate LTL properties from threat models to check their respective implementations’ Mealy machines.
-Also we can improve the learning process by using timing [64] and by extending the model with unknown inputs using input symbol mutation; if the unknown input triggers new behavior, we can dynamically extend the input alphabet (as described here [65]).
-This allows for a more holistic testing.
-11 Acknowledgement
-This work has received funding from the European Union under the Horizon Europe programme under grant agreement No.
-101168438 (project INTACT) as well as from the Swiss State Secretariat for Education, Research and Innovation and UK Research and Innovation under grant No.
-10126241.
-We further acknowledge the support of the Swedish Knowledge Foundation via the industrial doctoral school RELIANT under grant No.
-20220130.
-The authors want to thank the anonymous reviewers of the Computers & Security Journal, as well as of the International Conference on Runtime Verification for their valuable input.
-We further want to thank Thomas Grandits for helping to formulate requirements and properties more understandable.
+Listing 4: Example of the eMRTD property file.
+property {
+define {
+ERROR = system.error;
+EF = system.ef;
+DF = system.df;
+PROT = system.prot;
+AUTH = system.auth;
+S0 = (system.state ==0);
+S1 = (system.state ==1);
+S2 = (system.state ==2);
+S3 = (system.state ==3);
+S4 = (system.state ==4);
+S5 = (system.state ==5);
+UREADOK = system.ureadok;
+SREADOK = system.sreadok;
+READOK = system.readok;
+SSELEFOK = system.sselefok;
+SELEFOK = system.selefok;
+ACCESSOK = system.accessok;
+UACCESSOK = system.uaccessok;
+SACCESSOK = system.saccessok;
+INVKEYOK = system.invkeyok;
+WRONGKEYOK = system.wrongkeyok;
+OLDKEYOK = system.oldkeyok;
+CRIT = false;
+PRIV = false;
+Assertion{
+Assertion0: (!ERROR);
+LTL{
+P1:G( (!(!AUTH &&PROT)) !ACCESSOK);
+P2:G( (!PROT) (!UREADOK));
+P3:G( (!PRIV AUTH) && ((!(!PRIV &&CRIT)) !ACCESSOK));
+P4:G(!INVKEYOK);
+PlainRead: G(! (UREADOK && (!EF DF)));
+SecureRead: G(! (SREADOK && !(DF && AUTH && EF)));
+SecureSelect: G(! (SSELEFOK && !(DF && AUTH)));
+ReadFollowsSelect: U(!READOK,EF) G(!READOK);
+SecureReadFollowsSecureSelect: U(!SREADOK,SSELEFOK) G(!SREADOK);
+Appendix B UDS Rebeca Code Listings
+Listing 5: Example of generated eMRTD environment actor code.
+reactiveclass ENVIRONMENT(3) {
+knownrebecs {
+SYSTEM system;
+ENVIRONMENT() {
+self.start();
+void start() {
+system.req();
+msgsrv req_037f2700000000aa(){
+system.req();
+msgsrv req_065002003200c8aa(){
+system.req();
+Listing 6: Example of generated eMRTD system actor code.
+reactiveclass SYSTEM(3) {
+knownrebecs {
+ENVIRONMENT environment;
+statevars {
+boolean ext, auth, prog;
+int state;
+boolean error;
+boolean invkeyok, wrongkeyok, prot, accessok, uaccessok, crit, ureadok;
+msgsrv req() {
+error=false;
+invkeyok=false;
+wrongkeyok=false;
+prot=false;
+accessok=false;
+uaccessok=false;
+crit=false;
+ureadok=false;
+int data =?(0,1,2,[..],56);
+switch(data) {
+case 0: self.defaultsession(); break;
+case 1: self.programmingsession(); break;
+case 2: self.extendeddiagnosticsession(); break;
+case 3: self.communicationcontrol(); break;
+case 4: self.securityaccess(); break;
+case 5: self.securityaccesswithkey(); break;
+case 6: self.securityaccesswithwrongkey(); break;
+case 7: self.testerpresent(); break;
+case 8: self.checkaswbit(); break;
+sdefault: self.ERR();
+msgsrv securityaccesswithwrongkey(){
+if(state==0) {
+state=0; environment.req_037f277faaaaaaaa();
+} else
+if(state==1) {
+state=1; environment.req_037f2712aaaaaaaa();
+} else
+if(state==2) {
+state=2; environment.req_037f2724aaaaaaaa();
+} else
+if(state==3) {
+state=0; environment.ffffffffffffffff();
+} else
+if(state==4) {
+state=2; environment.req_037f2735aaaaaaaa();
+} else
+if(state==5) {
+state=5; environment.req_037f2724aaaaaaaa();
+} else
+if(state==6) {
+state=6; invkeyok=true;wrongkeyok=true;environment.req_02670aaaaaaaaaaa();
+} else
+if(state==7) {
+state=2; environment.req_037f2735aaaaaaaa();
+Listing 7: Example of the eMRTD property file.
+property {
+define {
+ERROR = system.error;
+EXT = system.ext;
+AUTH = system.auth;
+PROG = system.prog;
+S0 = (system.state ==0);
+S1 = (system.state ==1);
+S2 = (system.state ==2);
+S3 = (system.state ==3);
+S4 = (system.state ==4);
+S5 = (system.state ==5);
+S6 = (system.state ==6);
+S7 = (system.state ==7);
+INVKEYOK = system.invkeyok;
+WRONGKEYOK = system.wrongkeyok;
+PROT = system.prot;
+ACCESSOK = system.accessok;
+UACCESSOK = system.uaccessok;
+CRIT = system.crit;
+UREADOK = system.ureadok;
+PRIV=false;
+Assertion{
+Assertion0: (!ERROR);
+LTL{
+P1:G( (!(!AUTH &&PROT)) !ACCESSOK);
+P2:G( (!PROT) (!UREADOK));
+P3:G( (!PRIV AUTH) && ((!(!PRIV &&CRIT)) !ACCESSOK));
+P4:G(!INVKEYOK);